### v1

# ALMA-IMF XV: NH kinematic analysis on the intermediate protocluster G353.41 Key Words.: stars: formation – ISM: clouds – ISM: kinematics and dynamics – ISM: molecules1 Introduction

The ALMA-IMF Large Program provides multi-tracer observations of 15 Galactic massive protoclusters at matched sensitivity and spatial resolution. We focus on the dense gas kinematics of the G353.41 protocluster traced by NH (10), with a critical density of cm, and spatial resolution 0.02 pc. G353.41, at a distance of 2 kpc, is embedded in a larger scale (pc) filament and has a mass of M within pc. We extract the NH (10) isolated line component and we decompose it by fitting up to 3 Gaussian velocity components. This allows us to identify velocity structures that are either muddled or impossible to identify in the traditional position-velocity diagram. We identify multiple velocity gradients on large ( 1 pc) and small scales (0.2 pc). We find good agreement between the NH velocities and the previously reported DCN core velocities, suggesting that cores are kinematically coupled to the dense gas in which they form. We measure 9 converging “V-shaped” velocity gradients (km s pc) that are well-resolved (sizes pc), located in filaments, and are sometimes associated with cores near their point of convergence. The average timescales associated with the V-shapes are kyr, or about twice the free-fall time of cores in the same area (kyr) but substantially shorter than protostar lifetime estimates (0.5 Myr). We interpret these V-shapes as inflowing gas feeding the regions near cores (the immediate sites of star formation). We derive mass accretion rates in the range of M yr. This feeding might lead to further filament collapse and formation of new cores. We suggest that the protocluster is collapsing on large scales, but the velocity signature of collapse is slow compared to pure free-fall. Thus these data are consistent with a comparatively slow global protocluster contraction under gravity, and faster core formation within, suggesting the formation of multiple generations of stars over the protocluster lifetime.
ALMA-IMF XV: NH kinematic analysis on the intermediate protocluster G353.41
Key Words.:
stars: formation – ISM: clouds – ISM: kinematics and dynamics – ISM: molecules1 Introduction
While star clusters have been studied extensively over many decades at comparatively short wavelengths, their precursors, protoclusters have not been studied in depth until recently. Protoclusters (or embedded clusters) are the gas-dominated maternal environments where star clusters are born and whose stellar constituents will ultimately populate the field of our Galaxy. Protoclusters are distinct entities from star clusters. Both are defined as relatively compact configurations where the gravity is strong enough to influence the dynamics of their constituents. But in the latter, there is little to no gas, and the gravity of the cluster is dominated by the stars themselves. In protoclusters, in contrast, gravity is dominated by the cold gas in which the stars themselves are forming (Stutz & Gould, 2016; Csengeri et al., 2017; Stutz, 2018; Motte et al., 2018). Protoclusters are more accessible now than ever before thanks to ALMA and its exquisitely high resolution interferometric mm-wave data tracing the cold gas where the stars form (Sanhueza et al., 2019; Liu et al., 2020a; Motte et al., 2022). Inside protoclusters we witness the ongoing conversion of gas into compact and extremely dense stars, a process mediated by gas filaments (Stutz, 2018; González Lobos & Stutz, 2019; Álvarez-Gutiérrez et al., 2021) feeding gas structures called “cores” (André et al., 2010; Stutz & Kainulainen, 2015; Kuznetsova et al., 2015, 2018). Cores are compact gas mass concentrations, often defined to be of a size matching the resolution limit of the observations. In this case, we define cores to be order kau, for reasons described below.
In this paper, we focus on the G353.41 protocluster (see Fig. 1), and in particular, on the dense gas kinematics observationally accessible from the protocluster scale (2.9 pc) to the core scale. We trace this dense and cold gas using the NH (1-0) line observed with ALMA. Here NH, given its comparatively high critical density (see below), accesses the inner dense gas ”skeleton” of the protocluster structure, free from confusion induced by lower density gas. Meanwhile, ALMA permits us to obtain the resolution needed to trace structures down to the core scales where individual or small numbers of stars may be forming.
The ALMA-IMF Large Program111Proposal ID 2017.1.01355.L, PIs: Motte, Ginsburg, Louvet, Sanhueza (LP) maps 15 dense, nearby ( kpc), and massive ( 10 M) Milky Way protoclusters down to 2 kau scales (Motte et al., 2022), at matched spatial resolution. ALMA-IMF provides a large protocluster sample in order to test the universality of the stellar initial mass function (IMF) (Bastian et al., 2010; Offner et al., 2014). The ALMA-IMF LP also provides a vast catalogue of molecular lines, in bands 3 ( mm) and 6 ( mm). This rich molecular treasure trove allows for a detailed kinematical characterization of the gas, protostellar cores, and young stellar objects (YSOs) present in these protoclusters. The current publicly available ALMA-IMF data include, but are not limited to, continuum maps (Ginsburg et al., 2022; Díaz-González et al., 2023), 12 m data cubes of all spectral windows (Cunningham et al., 2023), core catalogues (Pouteau et al., 2022, Louvet et al. submitted), and hot core and outflow catalogues (Cunningham et al., 2023; Nony et al., 2023; Towner et al., 2024; Armante et al., 2024; Bonfand et al., 2024, Valeille-Manet et al. in prep). The data products derived from the ALMA-IMF LP allow us to constrain the different star forming environments, where we can analyze column densities, temperatures, outflow masses, core properties, and multi-tracer gas kinematics. This approach offers a thorough characterization of the processes taking place in these regions.
Field size | Pixel scale | Beam size | BPA | RMS | Channel width | RMS velocity range | V |
[°] | [K] | [ km s ] | [ km s ] | [ km s ] | |||
176″172″ | 0.72″ | 1.96″2.29″ | 80.19 | 0.37 | 0.23 | [43 ; 32], [0 ; +7] | 17 |
1.72 pc1.67 pc | 1.44 kau | 4 kau4.6 kau |
Motte et al. (2022) present a method of classifying these 15 protoclusters based on their evolutionary stage, assuming that they exhibit more H ii regions as they evolve. They take into account the flux ratio between the 1 mm to 3 mm continuum maps (S), and the free-free emission at the frequency of H41 (). They find that as protoclusters evolve, S decreases, while increases (Motte et al., 2022, see their Fig. 3). Using these constraints, they group their 15 protocluster as being in a young, intermediate, or evolved evolutionary state. Out of these 15 regions, we analyze the G353.41 protocluster (hereafter G353). In Fig. 1 we indicate the ALMA-IMF NH (10) coverage of G353 (centered at , (J2000) = 17:30:26.28,34:41:49.7) and its parent filament (dark lane traced by ATLASGAL 870 m emission; Schuller et al., 2009) with light green and gray contours respectively. Motte et al. (2022) classify this protocluster as being at an intermediate evolutionary state, located at 2 kpc, and hosting a total mass of 2.5 M. They describe G353 as isolated, without obvious interaction with massive nearby stellar clusters. Using moment maps derived from the NH (10) 12 m dataset they suggest the presence of multiple velocity components indicating a complex velocity field. They propose that G353 is composed of filaments interacting at the central hub. As presented in Bonfand et al. (2024), this region is an outlier in the ALMA-IMF hot core sample. Only one weak, low-mass ( 2 M) compact methyl formate source is detected and it lacks strong emission from complex organic molecules. They state that this protocluster is in a chemically poor stage, where further characterization of this region is required.
The NH (10) transition ( = 93.173809 GHz), given its high critical density, (Ungerechts et al., 1997), allows us to access the dense gas kinematics present in the innermost parts of star forming regions (Caselli et al., 2002a; Bergin et al., 2002; Tafalla et al., 2004; Lippok et al., 2013; Storm et al., 2014; Hacar et al., 2018; Chen et al., 2019; González Lobos & Stutz, 2019; Álvarez-Gutiérrez et al., 2021). The J = 10 transition presents seven hyperfine components (Cazzoli et al., 1985; Caselli et al., 1995, 2002a). The kinematic analysis of this complex emission can be simplified by considering only the well separated isolated component (93.17631 GHz; ; Cazzoli et al., 1985). Such simplification is convenient to study the complex velocity fields found at the center of filaments. These regions present the densest environments for star formation, usually presenting multiple, blended velocity components, where the velocity distributions exhibit twists, turns, spirals, and wave-like patterns (Csengeri et al., 2011; Fernández-López et al., 2014; Stutz & Gould, 2016; Liu et al., 2019; González Lobos & Stutz, 2019; Álvarez-Gutiérrez et al., 2021; Sanhueza et al., 2021; Redaelli et al., 2022; Olguin et al., 2023). Recent techniques, such as the intensity-weighted position-velocity (PV) diagrams (González Lobos & Stutz, 2019; Álvarez-Gutiérrez et al., 2021), allow us to characterize processes such as infall, outflow, or rotation present in these environments, where high spatial and spectral resolution studies open a window into the small scale gas kinematics of star forming regions. In addition to the PV diagrams, we can create Position-Position-Velocity (PPV) diagrams, in order to identify coherent structures that might be both spatially and kinematically associated (Chen et al., 2019; Henshaw et al., 2019; Sanhueza et al., 2021; Redaelli et al., 2022).
In this paper we investigate the NH dense gas kinematics of G353 from large (protocluster) to small (cores) scales. In § 2 we present the data. In § 3 we introduce our NH isolated extraction procedure. In § 4 we model and decompose the multiple velocity components found in the NH isolated component spectra. In § 5 we show our gas kinematic analysis, from protocluster to core scales. In § 6 we show that G353 might be under gravitational collapse at small and large scales. In § 7 we estimate mass accretion rates for multiple velocity gradients characterized in our NH data. We discuss our results in § 8, and we present our summary and conclusions in § 9.
2 Data
2.1 ALMA-IMF data
We make use of the NH (10) 12 m, 7 m, and Total Power observations described in Motte et al. (2022) for our analysis, providing robust uv plane coverage. We image the combination of the NH 7 m and 12 m (from now on called “7m+12m”) measurement set of G353, using the publicly available imaging scripts from the ALMA-IMF github repository333https://github.com/ALMA-IMF/reduction. These data are corrected by the primary beam response pattern. Due to the missing large-scale emission, we find that near the V of the protocluster (17 km s; Wienen et al., 2015; Motte et al., 2022) some subregions in the 7m+12m cube present deep negative artifacts (“negative bowls”). To cover all possible uv scales, we combine the NH 7m+12m continuum-subtracted cube with the Total Power observations from the ALMA-IMF LP. We use the feather444https://casa.nrao.edu/docs/taskref/feather-task.html task from CASA 5.6.0. With this combination, we were able to recover the missing flux, seen as negative bowls, present in the interferometric-only data. We produce a fully combined, multi-scale, feathered dataset which we use for our dense gas kinematic analysis.
To estimate and subtract the continuum emission present in the 7m+12m cube, we use the imcontsub555https://casa.nrao.edu/docs/taskref/imcontsub-task.html CASA task. We select the emission-free channels between 43 km s and 33 km s, and set the polynomial degree of the continuum fit (fitorder) to 0. We list relevant final image parameters in Table 1, such as the field size, pixel scale, beam size, root-mean-square noise (RMS), and channel width.
2.2 Core properties from published catalogues
We use the 1.3 mm continuum derived cores catalogue666Available at www.almaimf.com, extracted using the getsf algorithm on the 1.3 mm continuum maps, smoothed at a common resolution of 2700 au (45 sources, Louvet et al. submitted). We also use the DCN core velocities (15 sources, Cunningham et al., 2023) and the SiO outflow catalogue (16 sources, Towner et al., 2024) in order to look for correlation between the NH gas kinematics and cores/outflows position and properties. It is worth mentioning that, within a radius of from the center of G353 (Motte et al., 2022), we find 60% of the 1.3 mm cores (27 sources), and of the cores with DCN velocities and SiO outflows (11 sources from each catalogue). Of these 11 outflows, 7 are “red” 3 are “blue” (monopolar), and 1 is “bipolar” (Towner et al., 2024). The presence of these sources might imply a complex velocity field in this region, given that cores and outflows disturb the kinematics of the surrounding gas.
3 NH isolated component extraction
The NH (10) transition is characterized by its hyperfine emission composed by seven components (Caselli et al., 1995, see their Fig. 1). We present an ideal example of NH emission in Fig. 3, panel d. In this work we refer to the triplet of hyperfine components that present the highest intensities as the main NH components, located at the center of the line emission at GHz. We refer to the most blueshifted hyperfine component as the isolated component, at GHz, shifted by km s relative to the main NH component (see Table 1 from Cazzoli et al., 1985). We developed an algorithm to extract only the isolated hyperfine component from every pixel in the feathered datacube. This is in order to reduce the complexity of our data, given that it may contain multiple velocity components in addition to the hyperfine line emission. Considering that the NH emission moves in velocity across the protocluster, our approach is to find the velocity where the emission of the isolated component ends and remove the rest of the line emission. We also preserve the emission-free channels, at low ( km s to km s) and high ( km s to km s) velocities, to improve future RMS estimations if needed. Note that in the procedures described below, we use findpeaks777https://docs.scipy.org/doc/scipy/reference/generated/scipy.signal.find_peaks.html to detect peaks and valleys in different spectra.
Our extraction approach is separated into two procedures, for low and for high signal-to-noise (SNR) data. In order to determine which data have low or high SNR, we obtain the mean spectrum over all the spatial pixels of the cube, which serves as a guide to determine the velocity at the “mean dip” (V = -22 km s, dashed gray line in Fig. 3 panel “a”). This velocity represents the mean location of the intensity valley between the isolated and the main components of the NH emission. We define V = 3.2 km s as the difference between V and the velocity at the peak of the mean isolated component V (dashed black line in Fig. 3 panel “a”), used in our velocity guesses for the high SNR extraction procedure (see below).
To create a SNR map of the isolated component, we first measure the RMS noise in emission-free channels ( km s to km s), and the peak intensity in the channels range where the mean isolated component is located ( km s to V). This approach allows us to exclude the emission of the main line components. We encountered spurious emission at the edges of the SNR map. We adopt the procedure from Towner et al. (2024) by using the image processing techniques implemented by binaryerosion888https://docs.scipy.org/doc/scipy/reference/generated/scipy.ndimage.binary_erosion.html (1 iteration) and binarypropagation999https://docs.scipy.org/doc/scipy/reference/generated/scipy.ndimage.binary_propagation.html to clean the data for further analysis. binaryerosion allows us to remove the spurious emission in the outskirt of the map, although this approach also removes high SNR edges of our protocluster. Then, we use binarypropagation on the cleaned SNR map, using the original SNR map mask, to restore only the protocluster edges. To test our cleaning approach we compute the total integrated intensity using the Python package SpectralCube101010https://github.com/radio-astro-tools/spectral-cube in the range of 31.5 km s to V using the original and cleaned SNR mask. We estimate that the removed spurious emission accounts for 2% of the total integrated intensity for data with SNR.
In Fig. 2 we show the NH isolated component SNR map, where at SNR values 5 we capture the cloud emission while excluding noise (white contour). We set our isolated component SNR threshold to 5, in order to use one of the two extraction procedures (see below). In this section we refer to high (low) SNR spectrum if its isolated component (). For low SNR spectra, we extract all the channels in the velocity range from km s up until V (panel “b” in Fig. 3). For high SNR spectra the extraction procedure consists of creating different velocity guesses that represent the location of the intensity valley, similar to the definition of V (Fig. 3). Then, we select a velocity guess based on its associated weight (see description below). This approach is described in detail here:
First, we implement a rolling average along each spectra. This is in order to smooth over intensity bumps that might result in false positives for the detection of peaks and valleys. For this procedure, we average considering two channels before and after each velocity.
After smoothing, for each spectra we identify the isolated component peak using find_peaks. We call the velocity associated to this peak V. In the case of multiple velocity components it represents the most blueshifted one. We find the intensity valley between the isolated component and the NH main line emission by inverting the spectra and finding the first peak which is the inverted intensity valley. We define the associated velocity to this intensity valley as V.
We create three velocity guesses based on the properties of each spectrum in our cube (see points below). These are the 1 guess: VV. 2 guess: V, In the case of multiple isolated components this guess might incorrectly capture the intensity valley after the first isolated component. In that case, the other guesses are needed for a reliable isolated component extraction. 3 guess: VV to provide a velocity cut further away from the V. This guess is mainly useful in the case where multiple isolated components cover a velocity range larger than the one probed by the other two guesses.
From each velocity guess we estimate two parameters to later decide which one to use. One is the absolute value of its associated intensity “” (i.e. intensity at the guess velocity), and the other is distance in velocity “” to the mean dip. The subscript represents the guess associated to these parameters. We save the parameters of each guess in the lists “” and “”.
We normalize these lists by their minimum value ensuring that the guess with the smallest “” and “” will have a weight () of 1, defined in Eq 1. We do not encounter divergences in this normalization given these parameters are not exactly zero.
(1) where the “norm” subscript indicates that the parameter list is divided by its minimum value.
By visual inspection we consider that we obtain good extraction results when the weight is mostly dependent on and in a minor part on . This is reflected by the 0.2 and 0.8 factors multiplying and respectively, in the definition of “” in Eq 1.
We choose the guess with the weight closest to unity.
Similarly as for SNR 5, we extract the spectra from km s up until the velocity of the chosen guess, and preserving the emission-free channels from 0.7 km s to 6.7 km s.
Various examples of NH spectra and isolated hyperfine component extraction are shown in Fig. 3, where we can see spectra containing one (panel “b”), two (panel “c”), and three (panel “d”) velocity components, all well extracted by our procedure. In Stutz et al. (in prep) this approach is generalized to all ALMA-IMF regions for NH, providing reliable results. In Fig. 4 we show the moment 0 map estimated over the extracted NH isolated velocity components. We use the FilFinder Python package (Koch & Rosolowsky, 2015) in order to identify the most prominent filaments in this region (black lines in Fig. 4). We see that in the plane of the sky (POS) most of the 1.3 mm cores (red ellipses) are located on top of the filaments. This spatial agreement between filaments and protostelar cores is consistent with filamentary fragmentation (André et al., 2010; Busquet et al., 2013; Stutz & Kainulainen, 2015; Kuznetsova et al., 2015, 2018). The procedure, including the parameters we used for the filamentary identification, is presented in Appendix A.
4 NH isolated component velocity decomposition
In Fig. 3 we see that clear multiple isolated velocity components are present in our dataset. To characterize the complex dense-gas kinematics traced by NH we follow the method in Álvarez-Gutiérrez et al. (2021), and we use the spectroscopic toolkit PySpecKit (Ginsburg & Mirocha, 2011; Ginsburg et al., 2022) to model and decompose the isolated component emission. PySpecKit adjusts a fixed number of components set by the user, based on visual inspection of the data we impose three velocity components to every spectra and then remove false positives (see below). Given the kinematic complexity of the data and cursory inspection of the spectra, a simpler analysis with only two components contradicts the data. In essence, three components is the simplest possible choice, given the data. While this might fail for a small number of spectra that could require velocity components, the residuals indicate that this could occur in a severe minority of cases, and hence more components is not warranted given the SNR and resolution of this particular data set. To improve the convergence of PySpecKit, we create a set of ranges for the parameters that define each of the three Gaussian velocity components, namely the peak intensity, central velocity, and velocity dispersion. After testing different parameter ranges, we set the intensity range between 1.76 K (4 times the mean RMS) and 30 K, the velocity centroid from 30 km s to 20 km s, and the velocity dispersion from 0.22 km s to 1 km s.
From the results using the ranges defined above, we notice that some modeled components do not fit any emission. These fits are the result of imposing to the fitter a fixed number of components, given these spectra can be better represented by one or two velocity components. In these fits there is no uncertainty estimation for both the peak intensity and velocity dispersion. Based on these two criteria we remove those velocity fits from the modeled cube. With this cleaning approach we are left with spectra characterized by one (, two (, and three ( Gaussian velocity components. We present the Gaussian fits of the high SNR spectra from Fig. 3 in Fig. B.1.
In Fig. 5 we show the spatial distribution of the multiple Gaussian velocity components. In gray we indicate the main filamentary structure in the region (see § 3). The first and second velocity components, in blue and green respectively, present most of the high intensity emission and they also spatially dominate over the third, most red-shifted component. Both the first and second components trace mostly the filaments F1 and F3 from Fig. 4, where most of the 1.3 mm cores are located. The position of these cores coincide with high integrated intensity regions in these isolated velocity components. The most redshifted component is compact and less intense compared to the first and second velocity components. This velocity distribution is located mostly along the filament F2 and the central hub (see Fig. 4). In Fig. 6 we present the number of Gaussian velocity components for each spectra, where we highlight that:
Most of the NH data presents emission characterized by two velocity components.
Most of the spectra described by three velocity components are located in the innermost parts of the region.
Most of the cores (black ellipses; Louvet et al. submitted) are located in regions with spectra presenting two to three Gaussian velocity components, indicating kinematic complexity even at 4 kau (NH spatial resolution).
Single velocity component spectra are located preferentially in the outskirt of the protocluster.
In Fig. 7 we show the histogram of the fitted velocity centroid of each Gaussian velocity component. The peaks of these distributions are located at 27, 24.7, and 23.3 km s respectively, well-separated in velocity. From hereafter we refer to these distribution as blue, green, and red respectively. Most of the velocity components appear to be associated with the blue and green distributions. For consistency with the different tracers used in further analysis, we shift the isolated component velocities by +8 km s, to the reference frame of the main line components of NH (Cazzoli et al., 1985).
4.1 DCN & NH derived core velocities
In this section our goal is to increase the sample of core velocities from the already published DCN catalog, aiming to explore all the potential in these types of dataset. Given the relatively high of DCN (32) (cm) compared to NH (10) (), DCN (32) is known to coincide well with continuum peaks associated to cores (Liu et al., 2015; Cunningham et al., 2016; Minh et al., 2018), while NH is characterized by tracing the dense gas at the innermost parts of star forming regions (Fernández-López et al., 2014; Hacar et al., 2018; González Lobos & Stutz, 2019).
In Cunningham et al. (2023) they use ALMA-IMF 12 m observations of DCN (32) to study cores kinematics. They apply line emission fits for the DCN spectra inside the 1.3 mm cores from Louvet et al. (submitted). For this procedure they determine core velocities in all ALMA-IMF targets. They classify as DCN single and complex core velocities, spectra that can be fitted with one or multiple Gaussian velocity components respectively. Due to a global conservative SNR threshold the DCN fitting process missed the velocity estimation of some cores. For G353 only 15 out of the 45 cores present DCN velocity fits.
We use the ALMA-IMF DCN 12 m data from Cunningham et al. (2023), which presents a velocity resolution of km s. For each DCN velocity core described by a single component (Cunningham et al., 2023) we compare the emission of the DCN and modeled NH isolated spectra. We find an average velocity offset between the DCN peak and the closest NH isolated component peak of km s, less than two DCN channel widths. We use this approach for the remaining 30 cores, in order to determine their DCN velocities.
Here we estimate the RMS of the DCN data in emission-free channels in the range of 42 km s to 25 km s and we obtain the SNR map by dividing the peak intensity by the RMS. For the procedure below we only use DCN spectra with SNR 3. Next, we extract the average DCN and modeled NH isolated component spectra of these 30 cores. We identify the NH isolated velocity component closest to the DCN peak within three DCN channel widths. We find that 11 out of these 30 cores present DCN with SNR close to one NH velocity component. Here, we define the velocity of these cores as the velocity where the DCN emission peaks. On average, these cores have a velocity offset between these two tracers less than 0.8 km s ( DCN channels), similar to the results obtained for the 15 cores with DCN velocities from Cunningham et al. (2023), and they present an average velocity offset of 0.38 km s. Throughout this paper we refer to these cores as “DCN & NH cores” given they are derived from the comparison of these two tracers. In Table C.1 we include these DCN & NH core velocities, complementing the DCN catalogue from Cunningham et al. (2023). In Fig. C.1 we present two examples of DCN and NH spectra in cores where we see clear agreement between these tracers.
5 Analysis of position-velocity diagrams
5.1 Traditional PV diagram
We start by analyzing the “traditional” PV diagram shown in Fig. 8. We create this diagram by taking the total intensity along the Galactic longitude, where indicates the distance in parsec relative to the center of G353, assuming a distance to the protocluster of of 2 kpc (Motte et al., 2022). We see general agreement between the DCN core velocities and the NH velocity distribution. This suggests that most of the cores are still kinematically coupled to the dense gas in which they formed. As presented in § 4.1, the DCN and NH velocities match within 0.8 km s ( DCN channels).
[…]
Regarding the dense gas velocity distribution, in Fig. 8 we see a velocity spread of km s in the sub-region between 0.3 pc to 0.1 pc. Most of the intensity on this diagram is located at the upper part of this sub-region, at . This spread is also present in the PV diagram along and shown in the top right and bottom left panel of Fig. 9. We explore the possible origin of this structure in § 6.
5.2 Intensity-weighted PV diagrams
In the top left panel of Fig. 9 we show the spatial distribution of the fitted Gaussian velocity components (see § 4). The blue, green, and red color maps indicate the integrated intensity of the first, second, and third velocity components of the NH spectra respectively. Note that the spatial overlap between any of these components is presented in Fig. 6.
As seen in Fig. 8, the traditional PV diagram provides information on the dynamics on the large, protocluster-scale, environment. Meanwhile, the intensity-weighted position-velocity diagram (Fig. 9), where the color of each point indicates its integrated intensity, highlights the small core-scale kinematics. Similarly as in González Lobos & Stutz (2019) and Álvarez-Gutiérrez et al. (2021), from the isolated component line decomposition (§ 4), we derive the integrated intensity and velocity centroid for each Gaussian velocity component. Using these parameters we create intensity-weighted PV diagrams along the and coordinates. We present these NH PV diagrams in the bottom left and top right panels of Fig. 9. The key features on the position-position (PP) and on the top right PV diagram, are:
The agreement between the DCN core velocities and the overall NH PV structures suggests that cores are still kinematically coupled to the dense gas in which they formed.
We see at least nine clear and prominent V-shaped velocity gradients (see Fig. D.1), across all velocity components. The orientation of these V-shape, pointing to the left/right (top right panel) or up/down (bottom left panel), follow no clear preference.
In some cases, the vertex of these V-shapes is close spatially and in velocity to the location of cores.
In the plane of the sky (POS), all three velocity components overlap in most of the region.
This technique recovers the large scale velocity spread present in Fig. 8 and highlights small scale structures.
The most prominent V-shape is located at pc, 20.5 km s), between two 1.3 mm cores with DCN detections (see § 6).
For a better visualization of the 3D structure of these V-shapes we provide an interactive 3D PPV diagram at: rodrigoalvarez.space/research/figures.
5.3 Velocity gradients
In this section we focus on the most prominent blue V-shape (Fig. 9, top right panel). In Fig. 10 we show this velocity distribution in detail. In order to characterize the VGs composing this V-shape, we apply a linear fit to both the upper and lower VG. Given the visual linearity of the VGs composing the V-shape, we apply a linear fit to these distribution in order to characterize them. For these fits we consider data only above an integrated intensity threshold of 8 K km s and 3 K km s, for the upper and lower gradient respectively. We remove data not related to the velocity gradient, clustered in the ranges of , which lie just outside the filament hosting this V-shape on the POS. Additionally, we weight each point based on their integrated intensity to make our fits more robust. The slopes of the linear fits represent the VGs in km s pc. These linear fits follow the VGs distribution and these are somewhat asymmetric, the upper gradient is slightly shallower than the bottom gradient. Given the unknown inclination angle () of these structures relative to the POS, the observed VG is just a fraction of the original VG. These are related as VG = VG. These VGs present values between to km s pc (see Fig. 10). Additionally, we estimate the center of this V-shape as the velocity-weighted mean position of the points composing this structure. With this approach the position of the points closest to the V-shape apex present more weight, obtaining the center of this V-shape at () = (353.4135, 0.3657). This position is located between “core 2” and “core 3”, both of them having DCN velocity fits (Cunningham et al., 2023). These core present masses of 20.7 and 6.4 M respectively. We inspect the core catalog derived from the map at native resolution (Louvet et al. submitted) and the location of this V-shape do not coincide with any core.
In the left panel of Fig. 11 we show the integrated intensity of the multiple modeled NH isolated components. With colored boxes we show the areas where we create the different PV diagrams presented on the right panel. These boxes are centered at the main blue V-shape, matching the area of this V-shaped structure (see Fig. 12). We show that the overall structure in PV space is conserved at different angles, excluding the possibility of this velocity feature being the result of projection effects.
For this V-shaped structure, we note that the NH velocity range is V 2 km s, while the whole NH isolated component dataset present a V 8 km s (Fig. 9). From these VGs we derive timescales as = 1/VG, similar to the procedure for a rotating filament presented in Álvarez-Gutiérrez et al. (2021). We obtain timescales between kyr. These timescales are short compared to the 0.21 Myr free fall time (t) of the protocluster (Motte et al., 2022), and a few times larger than the t of nearby cores ( 20 kyr, within 0.1 pc of this V-shape). To determine the cores t, we use the 1.3 mm core masses from Louvet et al. (submitted). We characterized eight more NH V-shaped structure. These are presented in Appendix D. Note that within a beam size from the apex of V-shape “B” (see Fig. D.2), the continuum core “7” (M) is located.
In Henshaw et al. (2014), they propose two scenarios that might produce these V-shaped velocity gradients (see their Fig. 12). One scenario suggests that gas in a filament is flowing towards a denser region (infall), while the other scenario suggests that a protostellar outflow moves the dense gas located in its vicinity. To analyze the different dynamical processes present in this region, we use the ALMA-IMF 12 m data of the shock, outflow tracer SiO (54), from Cunningham et al. (2023). From those data we create its intensity-weighted PV diagram, presented in Fig. E.1 (see Appendix E for more details). We note that there is almost no SiO emission nor outflow sources at the location of the NH V-shape (see Fig. 12). The V within 10 from this main blue V-shape is 40 km s, while for the whole SiO dataset is 80 km s, showing a clear difference in the SiO V and the core velocities. Furthermore, the SiO V is times larger than the one of NH. This difference in traced velocities implies that these two molecules trace vastly different physical phenomena. We suggest that the small velocity range probed by NH indicates that the velocity gradients can be considered as infall signatures (see § 6). We discuss the possible morphology of the filaments hosting V-shapes in § 8.
6 G353 as a collapsing region
We use the SiO () and CO () data from Towner et al. (2024) and Cunningham et al. (2023) to identify possible outflows near the V-shape presented in Fig. 10. For CO we measure the RMS in the emission-free velocity range of 145 to 286 km s. We use only CO data with SNR3 for our analysis. The cleaning of the SiO data is described in Appendix E. In Fig. 12, with a fuchsia “” we indicate the center of the V-shape from § 5.3, located between two 1.3 mm cores 2 & 3 (black ellipses), which present DCN velocity detections. From these diagrams we see there is neither SiO nor CO outflow detection at the location of the main blue V-shape (, ) = (0 pc, 0 pc). In the right panel we show the position of the data composing this V-shape, where the velocity peaks towards the center of this velocity feature.
We derive the mass-weighted mean position of the two cores (2 & 3) closest to the V-shape to determine their barycenter (yellow “” in Fig. 12). We find that the difference between the center of the V-shape and the barycenter of these cores is ″(au), well below the beam size of the NH data. A similar offset is also present in the intensity and velocity profiles along filaments from ATOMS data (Zhou et al., 2022, see their Fig. 6). This small spatial offset might suggest that the gas flowing in the V-shape is produced by the gravitational pull towards the barycenter of cores 2 & 3, where the NH radial velocities peak. This interpretation is similar to the one proposed in Zhou et al. (2023) in their kinematic analysis of the G333 complex. They describe the V-shaped velocity gradients as the result of gas funneling from the molecular cloud to clumps which is then funneled into cores (see their Fig. 9) consistent with gravitational acceleration.
In Fig. 13 we show the mean spectra of different tracers at the central position of the main blue V-shape. These spectra are measured over a circular region with diameter equal to the major axis of the NH beam (2.28 ″; ). This circular area results in a coverage of 1.14 times the NH beam, and 5.6 times the beams of the HCO, DCN, and HCO data. We see NH and HCO present double component spectra with asymmetric peaks. Between these peaks we detect DCN and HCO emission. The asymmetric spectra present in NH and HCO is consistent with the “blue asymmetry” spectral feature, usually interpreted as infall signature, suggesting that this region is under gravitational collapse (e.g. Anglada et al., 1987; Mardones et al., 1997; Lee et al., 1999, 2001; Smith et al., 2012). Based on the idea that the V-shapes are the result of flowing gas along filaments towards denser regions, the blue-asymmetry detected at the center of the main blue V-shape suggests that gravitational collapse is taking place at the apex of the V-shaped structure.
Regarding large scales, in the traditional PV diagram presented § 5.1 (see Fig. 8), we see a clear velocity spread around pc, also present in the top right panel of Fig. 9. Below we compare this velocity spread with the velocity distribution produced by infall, where the gas velocities increase as the distance to the center of infall (“”) decreases:
For this comparison we model a sphere with a total mass of 150 M, a radius of 0.5 pc, and a power law density profile described by:
we provide the derivation of in Appendix F. was determined by visual inspection by comparing the obtained radial velocities of the model (see below), at different values, with the overall shape of the PV distribution.
We then estimate the infall velocity of each point, based on the cumulative mass distribution (“M”) at any given distance to the center (Eq. 2). We obtain the radial component of the infalling velocities as:
where represents the horizontal coordinate in the POS, while represents the (non-observed) depth of the sphere.
In Fig. 14, we show the coverage of the PV distribution from our model with a solid white line. We find good agreement between the PV distributions of our approach and the data. The PV distribution of our infall model is consistent with previous work that provide the expected PV distributions for spherical protostellar envelopes under infall (Tobin et al., 2012). At large scales we interpret the agreement between the PV diagrams of our model and the data as protocluster scale collapse due to gravitational contraction. It is worth noting that the mass of our model 5.5 times lower than the mass derived from the (H) map (Díaz-González et al., 2023). We speculate that a model considering complex processes such as turbulence, radiative transfer, and magnetic fields might solve this mass discrepancy while still matching the observed PV distribution.
7 Mass accretion rates in the V-shaped structure
Based on the idea that the V-shapes are a result of gas flowing toward cores, in this section we provide estimates of their mass accretion rates () for NH and H.
7.1 NH mass accretion rate
To derive the associated to the main blue V-shape (Fig. 10), we need to estimate its NH mass. For this procedure we use PySpecKit with the n2hp_vtau fitter, to fit the full NH line. The fitted parameters (see below) allow us to derive the (NH). The V-shape structure contains 77, 143, and 25 NH spectra with one, two, and three velocity components respectively. The bluest velocities in the three velocity component spectra accounts for 2% of the total number of velocities in this V-shape. For this reason, we model the full NH hyperfine line structure with one and two velocity components. In Table 2 we list the parameters and ranges used for this procedure.
After obtaining the modeled NH cube, we remove modeled components where , where and represent the estimated opacity and its associated error, respectively. This criterion is to ensure that we use reliable fitted parameters to determine our (NH) values. For the fitting of two NH components we only analyze the most blueshifted component. The resulting opacities follow a log-normal distribution with a peak at .
We derive the (NH) of the V-shape by using Eq. 5 (Caselli et al., 2002b):
Where , T, and are the opacity, excitation temperature, and velocity dispersion respectively, obtained from the full line fitting. The Planck and Boltzmann constants are represented by and respectively, and are the frequency and wavelength of NH, is the Einstein coefficient of the NH (10) transition, and are the statistical weights of the lower and upper energy levels, is the partition function, and is the energy of the lower level.
From the above procedure, inside the V-shaped structure, we get a total (NH) = 5.2410 cm and a total (NH) = 5.910 M. We use the average timescale of the VGs from § 5.3 (Fig. 10), = 64.5 kyr, to determine the (NH) as:
Note that the (NH) estimate (and (H) below) should be multiplied by , in order to account for the unknown inclination angle () of the protocluster/filaments relative to the POS.
Excitation temperature [K] | (T) | 2.73 80 |
Opacity | () | 0.01 40 |
Centroid velocity [km s] | () | 25 15 |
Velocity dispersion [km s] | () | 0.20 3 |
7.2 H mass accretion rate
To ensure that we estimate the V-shape (H) on the same area as in the NH hyperfine line fitting, we use the CASA task imregrid to obtain the continuum-derived (H) map from Díaz-González et al. (2023) at the resolution of the NH data. We determine that in this V-shape the total (H 1.03 cm. Here, we derive a (H) map using Eq. 7:
where from this (H) map we consider only the points that are part of the V-shape. To determine the mass associated to flowing motions we subtract the core masses (from Louvet et al. submitted) that are located inside this V-shape. Note here that this mass map is an upper limit given that we do not apply a background correction. We obtain a total of (H 54 M. Considering used in Eq. 6, we derive the (H) as:
We use the procedure described in this section to estimate the (H) of other eight V-shapes shown in Fig. D.2. We include these values in Table D.1. The average (H) of these V-shapes is 3.4 10 M yr. We estimate the free-fall time of all 45 1.3 mm cores from Louvet et al. (submitted). These values present large scattering, ranging from M yr, with 28 of these cores presenting (H) M yr. For cores 2 & 3, the average (H) is 15.5 M yr, about twice the (H) of the main blue V-shape, located between these two cores.
7.3 NH relative abundance
For the main blue V-shape (Fig. D), we derive the NH relative abundance (NH), using the NH and H column densities obtained above, as:
The NH) value obtained above is consistent, within the scatter, with other estimates in massive Galactic star formation regions in the ranges of (1.6 - 3.8) 10 (Caselli et al., 2002a; Henshaw et al., 2014, Sandoval-Garrido et al. in prep.).
8 Discussion
The V-shaped velocity gradients described in this work have been detected across multiple Galactic star forming system. Stutz & Gould (2016) introduced the Slingshot mechanism in the Integral Shaped Filament (ISF) located in Orion A. They show undulations of the region in both position and velocity, suggesting that these features appear to be ejecting protostars (see their Fig. 2). Furthermore, Stutz (2018) characterize a standing wave in the neighborhood of the ISF, consistent with the Slingshot mechanism. It is possible that the undulations in the works above might result in the observed V-shaped structures seen in different studies (see below). González Lobos & Stutz (2019) identify six evenly spaced (every 0.44 pc) velocity peaks along the spine of the ISF in Orion A. They suggest that this periodicity is consistent with the wave-like perturbation in the gas caused by the Slingshot mechanism. In Álvarez-Gutiérrez et al. (2021) they analyze the L1482 filament located in the California Molecular Cloud. In all of the analyzed tracers there is a clear velocity peak in their north region (length 1.8 pc, mass M). This subregion contains a higher gas density and higher number of YSOs compared to the more quiescent south part.
While the two regions described above are considered nearby (both at D 500 pc), more distant regions also present these velocity features which we list below. Zhou et al. (2022) study the velocity profiles along filaments from the ATOMS survey (Liu et al., 2020b). The median mass of their sources is 1.4 M with a median length of the filaments at pc. By analyzing the HCO (10) emission they find converging VGs along filaments (see their Figs. 6 & 10), which they also detected using simulations from Gómez & Vázquez-Semadeni (2014). These VGs at scales comparables to the V-shapes presented here are consistent with our VGs estimates (see their Fig. 7 & 8). In Zhou et al. (2023) they analyzed CO (21) APEX/LAsMA data of the G333 complex. They identify multiple V-shaped VG (see their Fig. 7) which they describe as the PV projection of a funneling structure in PPV space (see their Fig. 9). The origin of this structure is due to material inflowing towards the central hub and also due to gravitational contraction of star-forming clouds or clumps.
Redaelli et al. (2022) use ALMA NH (10) isolated component data of the high-mass (5200 M) clump AGAL014.492-00.139 identifying multiple coherent structures in PPV space (trees “B” and “G”; right panel of their Fig. 7 & 9). These are characterized by multiple undulations, and possible V-shaped VGs. For their “G” PV distribution, they suggest that one scenario is where the dense gas is flowing along the filament (of length 0.2 pc) from protostar “p3” towards the protostar “p2”. This motion has an = 2.2 M yr, being in the range of the we derive for our VGs (see Table D.1).
In Rawat et al. (2024) they analyze CO(10) ALMA data observing a V-shaped VG (see their Fig. 14) along the ridge of the G148.24+00.41 (G148) cloud. This V-shape peaks towards the dense clump at the center of this region, possibly indicating gas inflow along their filaments F2 and F6 towards the hub. Note that the length of the V-shape in G148 is pc, while our most prominent V-shape (Fig. 10) span and lengths of their identified filaments are 1.3 to 6.9 M and 14 to 38 pc respectively, large compared to the total mass (2.5 M) and extent of G353 ( pc). This different in probed length and masses might be reflected by their mean VG km s pc, orders of magnitude smaller than our VGs.
In Pan et al. (2024) using APEX CO (21) data of the filamentary cloud G034.43+00.24 (G34) they identify converging VGs of lengths pc towards the “middle ridge” see their Fig. 3, top panel). They interpret these VG as gas flowing from the filaments onto dense clumps, located at the center of G34. These VGs of their southern and northern filaments are in the range of km s pc, and they estimate the total mass inflow rate towards the middle ridge as M yr, similar to our estimates.
Current work by Sandoval-Garrido et al. (in prep.) in G351.77 (intermediate protocluster, located at 2 kpc; Motte et al., 2022; Reyes-Reyes et al., 2024) use a similar analysis as we present in this work, where they identify multiple V-shaped velocity structures. In Salinas et al. (in prep) they analyze the kinematics of the evolved protocluster G012.80 (located at 2.4 kpc; Motte et al., 2022), where they implement similar techniques and find velocity signatures of filamentary rotation.
As presented here, V-shaped VGs appear to be a generic feature across a wide range of star forming environments, probing VGs with differences of up to orders of magnitude in spatial scales ranging from 0.1 to 10 pc. Based on these works, the V-shaped VGs appear to be a generic feature in star forming regions. Despite being commonly detected in recent studies, it is still not clear how they are produced. Henshaw et al. (2014) highlights the degeneracy regarding the opposite interpretations of these V-shaped velocity structures. They suggest that these VGs can be a signature of gas flows along kinked filaments towards a core located at their convergence point. From our analysis regarding the most prominent V-shape (see § 5.3 & 6) we see that no core is located at its apex, although cores 2 & 3 are within pc. Also the position difference between the center of the V-shape with the barycenter between these two cores is AU. This is consistent with the idea of small-scale gravitational collapse within the protocluster, similar to clump decoupling from their parent molecular cloud (Peretto et al., 2023). Based on this, we conclude that cores may be located in the vicinity of the velocity apex, and not necessarily on top of it. These gas flows towards denser regions may result in the formation of high-mass cores in later stages during the evolution of the protocluster.
Regarding the kinked morphology of the regions hosting V-shapes, one scenario regarding magnetized shocks is presented in Inoue & Fukui (2013) and Inoue et al. (2018). They use magnetohydrodynamics simulations to characterize the interaction of molecular clouds and a magnetized shock produced by a cloud-cloud collision. They find that the shock layer decelerates as it collides with denser regions. This deceleration reshapes the shock layer to be oblique, leading to the formation of kinked filaments and converging flows, which are oriented towards the apex of these filaments. They predict that magnetic fields present in the region should be perpendicular to these filaments and bend with the shock around the filament (Inoue et al., 2018, see their Fig. 3). In Bonne et al. (2020) and Bonne et al. (2023) they propose that this scenario takes place in the Musca and the DR21 filaments. In both of these regions they detect V-shaped VGs which they suggest are the result of cloud-cloud collisions bending the magnetic field (Bonne et al., 2020, see their Fig. 22 & 23). Further observations of magnetic field polarization in the POS, along with information along the line of sight is required to evaluate these models.
Another possibility is that these kinked structures could be caused by mechanisms such as the Slingshot. This mechanism proposes a standing wave, longitudinal gravitational instabilities, or large scale oscillations possibly caused by a possibly helical magnetic field morphology, causing ejections of protostars and protoclusters from their maternal filament (Stutz & Gould, 2016; Stutz, 2018; Stutz et al., 2018; Liu et al., 2019).
A different interpretation is that they are the product of out-flowing material coming from a forming protostar interacting with the surrounding dense gas (see their Fig. 12). To shed light into this degeneracy in G353 we compare the NH (dense gas tracer) radial velocities and SiO (shock/outflow tracer) as a proxy for energies. The velocity range V covered by the NH emission is 8 km s, while for SiO is 80 km s. Given the difference in probed velocities between SiO and NH (see § 5.3) and the analysis presented in § 6 we suggest that the V-shapes in G353 are a signature of infall.
As presented in this section, the V-shaped VGs appear to be common in recent studies of protoclusters. The multiple VGs that conform the V-shapes present in G353 have values of to km s pc, with timescales ranging from to kyr, and values . These values are similar to VGs in other regions with comparable sizes and masses. In G353 it is likely that these VGs are the result of dense gas moving through filaments, possibly increasing the density of the central regions, shaping the overall velocity field at large and small scales, and leading to a further increase of the core population and star formation activity.
One important aspect of the V-shapes that is still not well understood is the timescale associated of the VGs (VG). It is not clear if nor how these timescales determine core formation lifetimes or impact the star formation environment in general. In our sample of V-shapes the timescales are in tens of kyrs with the average value of kyr, times the cores , while the of the whole protocluster is Myr. In Rawat et al. (2024) they estimate the longitudinal collapse timescales for their filaments, being in the range of 5 15 Myr. Using their derived VGs we estimate their associated timescales to be between and 50 Myr, orders of magnitude larger than our small scale V-shapes timescales. We suggest that the VG timescales might serve as an upper limit for filamentary collapse timescales. In Zhou et al. (2022) they determine gas accretion times as a function of the lengths of their filaments, assuming that the VGs produced by gas inflow (see their Fig. 11). At filament lengths comparable to our V-shapes ( pc) their gas accretion timescales are on the order of our estimates (see Table D.1).
It is also interesting to consider the mass accretion rates measured here compared to the available protocluster mass reservoir to explore implications for the duration of the gas dominated phase. The total mass accretion rate of our V-shaped structures is 310 M yr (see Table D.1). Considering the total mass (M) of G353 as a mass reservoir, we estimate the time needed fully consume the gas (). Here we assume that is representative of flows feeding gas onto cores. We estimate M, where the total mass of the region is 2500 M (Motte et al., 2022). We obtain a = 0.8 Myr, of similar magnitude but about four times larger than the of the protocluster. Considering that our estimate of is certainly a lower limit (see discussion above), the actual value value of is likely to be shorter, so closer to the free-fall time estimate. Given that the protocluster does not appear to be in a state of free-fall (see § 6) but instead undergoing comparatively slow gravitational contraction, the similarity in these relatively crude estimates seems remarkable. While we do not yet have an explanation for why this relatively good match in timescales, it would seem to indicate that protocluster evolution may be a self-regulating process. Larger samples and similar analysis will test this hypothesis.
Moreover, the approximate concordance of and may indicate that the “phase transition” of protocluster gas mass being converted into stellar mass could contribute a relevant “negative pressure” counteracting effects of e.g. feedback over the lifetime of the protocluster.
9 Summary & conclusions
We characterize the complex dense gas kinematics of G353 using ALMA-IMF LP observations. The data used in this paper mainly consist of the fully combined NH data cube, but we also include 1.3 mm continuum cores and DCN cores velocity catalogues, SiO 12 m observation, and a (H) 1.3 mm continuum derived map. We summarize our main results below.
1.
With our NH isolated component modeling, we find that most of the 1.3 mm cores are located in regions with 2 to 3 velocity components. This indicates kinematic complexity down to 4 kau scales.
2.
We increase the number of cores with a V estimate in this region by further examining the DCN emission and comparing it with the NH data extracted towards the core positions. We find that 11 cores, which were previously undetected in the in the DCN background-subtracted fitting from Cunningham et al. (2023), are identified with our method. With this approach we increase our core velocities sample from 15 to 26, accounting for of the total 45 1.3 mm continuum cores. These are presented in Table C.1.
- 3.
4.
From the PV diagrams, we see the DCN core velocities are in agreement with the NH velocity distribution (within a few DCN channel widths). This suggests coupling between the cores and the dense gas in which they formed.
5.
In the intensity-weighted PVs we see clear V-shaped velocity structures, composed by two linear velocity gradients (VGs) converging into a common point. These VGs are present across all NH velocity components. Some of them are near the location of cores in both position and velocity (see § 5.2)
6.
We successfully characterize nine V-shaped VGs well detected in our NH data (see Fig. D.2).
- 7.
8.
We estimate the barycenter of cores 2 & 3, presenting an offset relative to the center of the V-shape of ″( 600 au) well below the beam size of our NH data.
9.
For V-shape “B” we find that core “7”, with a mass of 6 M, is located within a beam size from its apex.
10.
We suggest that the dense gas is flowing along the filament, producing the V-shaped structure towards the derived barycenter.
11.
We characterize the VGs composing our sample of V-shapes by applying linear fits to these distributions. We estimate timescales associated to the VGs as VG. These timescales are between 35 to 170 kyr, with an average of 67 kyr. These values are short compared to the t of the protocluster ( Myr), and times larger that the cores average t ( kyr).
12.
We suggest that at small scales the NH V-shaped structures indicate gas motions along filaments, towards denser regions.
13.
Using an H mass map and the V-shapes mean timescales, we derive H mass accretion rates of , consistent with previous studies on regions that present gas flows along filaments towards denser object or regions, such as protostars and clumps.
14.
In SiO, the PV structure covers a velocity range (V) of km s, while for NH V is km s. This difference suggests that NH is tracing infall, a less energetic processes compared to SiO, a shock and outflow tracer.
15.
We model a gravitationally collapsing sphere. The derived radial velocities are consistent with the large scale morphology of the traditional PV diagram. This agreement suggests that at large scales the G353 protocluster is undergoing gravitational contraction.
Overall, it is imperative to replicate the kinematic analysis presented in this work in the remaining ALMA-IMF fields and other Galactic star forming regions. By increasing the sample of analyzed fields we might find correlations between evolutionary state (young, intermediate, or evolved; see Motte et al., 2022), star formation activity, cores and outflow population properties, and their velocity field. This approach will allow us to better describe the kinematic processes taking place in this early stage of star formation.
Acknowledgements.
This paper makes use of the following ALMA data: ADS/JAO.ALMA#2017.1.01355.L. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. We thank Elena Redaelli and Diego R. Matus Carrillo for very helpful discussions. R.A. gratefully acknowledges support from ANID Beca Doctorado Nacional 21200897. A.S. gratefully acknowledges support by the Fondecyt Regular (project code 1220610) and ANID BASAL project FB210003. F.L. acknowledges the support of the Marie Curie Action of the European Union (project MagiKStar, Grant agreement number 841276) F.M. acknowledges support from the French Agence Nationale de la Recherche (ANR) under reference ANR-20-CE31-009, of the Programme National de Physique Stellaire and Physique et Chimie du Milieu Interstellaire (PNPS and PCMI) of CNRS/INSU (with INC/INP/IN2P3). R.G.M. and D.D.G. acknowledge support from UNAM-PAPIIT project IN108822 and from CONACyT Ciencia de Frontera project ID 86372. F.M., F.L., and N.C. acknowledge support from the European Research Council (ERC) via the ERC Synergy Grant ECOGAL (grant 855130). N.C. acknowledges funding from the ERC under the European Union’s Horizon 2020 research. P.S. was partially supported by a Grant-in-Aid for Scientific Research (KAKENHI Number JP22H01271 and JP23H01221) of JSPS. M.B. is a postdoctoral fellow in the University of Virginia’s VICO collaboration and is funded by grants from the NASA Astrophysics Theory Program (grant number 80NSSC18K0558) and the NSF Astronomy & Astrophysics program (grant number 2206516). T.Cs. has received financial support from the French State in the framework of the IdEx Université de Bordeaux Investments for the future Program. S.D.R. acknowledges the funding and support of ANID-Subdirección de Capital Humano Magíster/Nacional/2021-22211000. T.B. acknowledges the support from S. N. Bose National Centre for Basic Sciences under the Department of Science and Technology, Govt. of India. G.B. acknowledges financial support from the grants PID2020-117710GB-I00 and CEX2019-000918 funded by MCIN/AEI/10.13039/501100011033. A.K. and L.B. gratefully acknowledge support from ANID BASAL project FB210003. F.O. acknowledge the support of the Ministry of Science and Technology of Taiwan, projects No. 109-2112-M-007-008-, 110-2112-M-007-023-, and 110-2112-M-007-034-.References
- Álvarez-Gutiérrez et al. (2021) Álvarez-Gutiérrez, R. H., Stutz, A. M., Law, C. Y., et al. 2021, ApJ, 908, 86
- André et al. (2010) André, P., Men’shchikov, A., Bontemps, S., et al. 2010, A&A, 518, L102
- Anglada et al. (1987) Anglada, G., Rodriguez, L. F., Canto, J., Estalella, R., & Lopez, R. 1987, A&A, 186, 280
- Armante et al. (2024) Armante, M., Gusdorf, A., Louvet, F., et al. 2024, arXiv e-prints, arXiv:2401.09203
- Bastian et al. (2010) Bastian, N., Covey, K. R., & Meyer, M. R. 2010, ARA&A, 48, 339
- Bergin et al. (2002) Bergin, E. A., Alves, J., Huard, T., & Lada, C. J. 2002, ApJ, 570, L101
- Bonfand et al. (2024) Bonfand, M., Csengeri, T., Bontemps, S., et al. 2024, arXiv e-prints, arXiv:2402.15023
- Bonne et al. (2020) Bonne, L., Bontemps, S., Schneider, N., et al. 2020, A&A, 644, A27
- Bonne et al. (2023) Bonne, L., Bontemps, S., Schneider, N., et al. 2023, ApJ, 951, 39
- Busquet et al. (2013) Busquet, G., Zhang, Q., Palau, A., et al. 2013, ApJ, 764, L26
- Caselli et al. (2002a) Caselli, P., Benson, P. J., Myers, P. C., & Tafalla, M. 2002a, ApJ, 572, 238
- Caselli et al. (1995) Caselli, P., Myers, P. C., & Thaddeus, P. 1995, ApJ, 455, L77
- Caselli et al. (2002b) Caselli, P., Walmsley, C. M., Zucconi, A., et al. 2002b, ApJ, 565, 344
- Cazzoli et al. (1985) Cazzoli, G., Corbelli, G., Degli Esposti, C., & Favero, P. 1985, Chemical Physics Letters, 118, 164
- Chen et al. (2019) Chen, H.-R. V., Zhang, Q., Wright, M. C. H., et al. 2019, ApJ, 875, 24
- Csengeri et al. (2011) Csengeri, T., Bontemps, S., Schneider, N., Motte, F., & Dib, S. 2011, A&A, 527, A135
- Csengeri et al. (2017) Csengeri, T., Bontemps, S., Wyrowski, F., et al. 2017, A&A, 601, A60
- Cunningham et al. (2023) Cunningham, N., Ginsburg, A., Galván-Madrid, R., et al. 2023, A&A, 678, A194
- Cunningham et al. (2016) Cunningham, N., Lumsden, S. L., Cyganowski, C. J., Maud, L. T., & Purcell, C. 2016, MNRAS, 458, 1742
- Díaz-González et al. (2023) Díaz-González, D. J., Galván-Madrid, R., Ginsburg, A., et al. 2023, ApJS, 269, 55
- Fernández-López et al. (2014) Fernández-López, M., Arce, H. G., Looney, L., et al. 2014, ApJ, 790, L19
- Ginsburg & Mirocha (2011) Ginsburg, A. & Mirocha, J. 2011, PySpecKit: Python Spectroscopic Toolkit, Ver. 0.1.23, Astrophysics Source Code Library
- Ginsburg et al. (2022) Ginsburg, A., Sokolov, V., de Val-Borro, M., et al. 2022, AJ, 163, 291
- Gómez & Vázquez-Semadeni (2014) Gómez, G. C. & Vázquez-Semadeni, E. 2014, ApJ, 791, 124
- González Lobos & Stutz (2019) González Lobos, V. & Stutz, A. M. 2019, MNRAS, 489, 4771
- Hacar et al. (2018) Hacar, A., Tafalla, M., Forbrich, J., et al. 2018, A&A, 610, A77
- Henshaw et al. (2014) Henshaw, J. D., Caselli, P., Fontani, F., Jiménez-Serra, I., & Tan, J. C. 2014, MNRAS, 440, 2860
- Henshaw et al. (2019) Henshaw, J. D., Ginsburg, A., Haworth, T. J., et al. 2019, MNRAS, 485, 2457
- Inoue & Fukui (2013) Inoue, T. & Fukui, Y. 2013, ApJ, 774, L31
- Inoue et al. (2018) Inoue, T., Hennebelle, P., Fukui, Y., et al. 2018, PASJ, 70, S53
- Koch & Rosolowsky (2015) Koch, E. W. & Rosolowsky, E. W. 2015, MNRAS, 452, 3435
- Kuznetsova et al. (2015) Kuznetsova, A., Hartmann, L., & Ballesteros-Paredes, J. 2015, ApJ, 815, 27
- Kuznetsova et al. (2018) Kuznetsova, A., Hartmann, L., & Ballesteros-Paredes, J. 2018, MNRAS, 473, 2372
- Lee et al. (1999) Lee, C. W., Myers, P. C., & Tafalla, M. 1999, ApJ, 526, 788
- Lee et al. (2001) Lee, C. W., Myers, P. C., & Tafalla, M. 2001, ApJS, 136, 703
- Lippok et al. (2013) Lippok, N., Launhardt, R., Semenov, D., et al. 2013, A&A, 560, A41
- Liu et al. (2015) Liu, H. B., Galván-Madrid, R., Jiménez-Serra, I., et al. 2015, ApJ, 804, 37
- Liu et al. (2019) Liu, H.-L., Stutz, A., & Yuan, J.-H. 2019, MNRAS, 487, 1259
- Liu et al. (2020a) Liu, T., Evans, N. J., Kim, K.-T., et al. 2020a, MNRAS, 496, 2790
- Liu et al. (2020b) Liu, T., Evans, N. J., Kim, K.-T., et al. 2020b, MNRAS, 496, 2790
- Mardones et al. (1997) Mardones, D., Myers, P. C., Tafalla, M., et al. 1997, ApJ, 489, 719
- Minh et al. (2018) Minh, Y. C., Liu, H. B., Galvań-Madrid, R., et al. 2018, ApJ, 864, 102
- Motte et al. (2022) Motte, F., Bontemps, S., Csengeri, T., et al. 2022, A&A, 662, A8
- Motte et al. (2018) Motte, F., Bontemps, S., & Louvet, F. 2018, ARA&A, 56, 41
- Nony et al. (2023) Nony, T., Galván-Madrid, R., Motte, F., et al. 2023, A&A, 674, A75
- Offner et al. (2014) Offner, S. S. R., Clark, P. C., Hennebelle, P., et al. 2014, in Protostars and Planets VI, ed. H. Beuther, R. S. Klessen, C. P. Dullemond, & T. Henning, 53–75
- Olguin et al. (2023) Olguin, F. A., Sanhueza, P., Chen, H.-R. V., et al. 2023, ApJ, 959, L31
- Pan et al. (2024) Pan, S., Liu, H.-L., & Qin, S.-L. 2024, ApJ, 960, 76
- Peretto et al. (2023) Peretto, N., Rigby, A. J., Louvet, F., et al. 2023, MNRAS, 525, 2935
- Pouteau et al. (2022) Pouteau, Y., Motte, F., Nony, T., et al. 2022, arXiv e-prints, arXiv:2212.09307
- Rawat et al. (2024) Rawat, V., Samal, M. R., Walker, D. L., et al. 2024, MNRAS[arXiv:2401.03202]
- Redaelli et al. (2022) Redaelli, E., Bovino, S., Sanhueza, P., et al. 2022, ApJ, 936, 169
- Reyes-Reyes et al. (2024) Reyes-Reyes, S. D., Stutz, A. M., Megeath, S. T., et al. 2024, MNRAS[arXiv:2403.02456]
- Sanhueza et al. (2019) Sanhueza, P., Contreras, Y., Wu, B., et al. 2019, ApJ, 886, 102
- Sanhueza et al. (2021) Sanhueza, P., Girart, J. M., Padovani, M., et al. 2021, ApJ, 915, L10
- Schuller et al. (2009) Schuller, F., Menten, K. M., Contreras, Y., et al. 2009, A&A, 504, 415
- Smith et al. (2012) Smith, R. J., Shetty, R., Stutz, A. M., & Klessen, R. S. 2012, ApJ, 750, 64
- Storm et al. (2014) Storm, S., Mundy, L. G., Fernández-López, M., et al. 2014, ApJ, 794, 165
- Stutz (2018) Stutz, A. M. 2018, MNRAS, 473, 4890
- Stutz et al. (2018) Stutz, A. M., Gonzalez-Lobos, V. I., & Gould, A. 2018, arXiv e-prints, arXiv:1807.11496
- Stutz & Gould (2016) Stutz, A. M. & Gould, A. 2016, A&A, 590, A2
- Stutz & Kainulainen (2015) Stutz, A. M. & Kainulainen, J. 2015, A&A, 577, L6
- Tafalla et al. (2004) Tafalla, M., Myers, P. C., Caselli, P., & Walmsley, C. M. 2004, A&A, 416, 191
- Tobin et al. (2012) Tobin, J. J., Hartmann, L., Bergin, E., et al. 2012, ApJ, 748, 16
- Towner et al. (2024) Towner, A. P. M., Ginsburg, A., Dell’Ova, P., et al. 2024, ApJ, 960, 48
- Ungerechts et al. (1997) Ungerechts, H., Bergin, E. A., Goldsmith, P. F., et al. 1997, ApJ, 482, 245
- Wienen et al. (2015) Wienen, M., Wyrowski, F., Menten, K. M., et al. 2015, A&A, 579, A91
- Zhou et al. (2022) Zhou, J.-W., Liu, T., Evans, N. J., et al. 2022, MNRAS, 514, 6038
- Zhou et al. (2023) Zhou, J. W., Wyrowski, F., Neupane, S., et al. 2023, A&A, 676, A69
Appendix A Filamentary identification with FilFinder
Here we describe the procedure to identify the main filamentary structures presented in § 3 using FilFinder (Koch & Rosolowsky 2015).
For this approach we use the moment 0 map of the extracted NH isolated components that present a SNR 5. To estimate the moment 0 map we used the moment task from the SpectralCube Python package, within the velocity range of km s to 0 km s. As part of the pre-processing of the moment 0 map before the filamentary detection, we decrease the contrast in the image by using the preprocess_image task and its argument flatten_percent set to 90. Now, in order to indicate to FilFinder the area we to identify filaments use the subtask create_mask with the following parameters: glob_thresh: 4.5 K km s, size_thresh: 0.25 pc, smooth_size: 0.12 pc, border_masking: False, fill_hole_size: 0.013 pc. The resulting mask is presented in Fig. A.1 with a white contour.
Then, we obtain the skeletons of the mask by using medskel. The derived structures are presented with red and black lines in Fig. A.1. Given we are only interested in the large scale filaments, we use analyze_skeletons in order to “prune” the small scale structures. For this pruning we use branch_thresh: 0.3 pc, prune_criteria: ’length’, max_prune_iter: 0. This approach results in removing the small filaments (red lines in Fig. A.1) from the original skeleton and to obtain the main filamentary structure in G353 (black lines in Fig. A.1).
Appendix B Examples of the isolated components fitting
Appendix C DCN & NH derived core velocity
Appendix D V-shaped structures
In § 5.3 we characterized the most prominent V-shaped structure we detect in Fig. 9. We repeat this process for other eight different V-shaped structures, including the linear fits to the velocity gradients. In Fig. D.1 we indicate their location in PV space with arrows and present individual close-up for each V-shape in Fig. D.2. In Table D.1 we list their VGs, timescales, H masses, and mass accretion rates.
Here we list a few clarifications due to projection effects seen in these V-shaped structures:
In position-position space, only V-shape “B” presents a core within a beam size from its apex.
For V-shape “A”, the 1.3 mm core with DCN single velocity component, located at the apex of this V-shape, is not spatially related to it.
In V-shapes “G” and “H” we see the same 1.3 mm cores with NH velocities. These V-shapes are not the same distribution. They are overlapped in PV space and spatially separated by .
We improve the clarity of V-shape “B” by rotating the data in PP space by 80 counter-clockwise. We apply this process for V-shapes “E”, “G”, and “H” with an angle of 33 clockwise.
V-shapes G and H overlap in PV space but these are structures spatially separated.
Appendix E SiO Intensity-weighted position-velocity diagram
To create the SiO intensity-weighted PV diagram, first, we remove most of the noisy spectra by considering data with . Then, we estimate the integrated intensity and velocity centroid at each pixel. We find improvements in our cleaning by using only spectra with integrated intensity K km s. Using the coordinate, integrated-intensity, and velocity centroid of each spectrum, we create the SiO intensity-weighted PV diagrams we show in E.1.
Core | RA | DEC | F | F | PA | Mass | V | Type |
Number | [ ] | [ ] | [″] | [″] | [ ] | [M ] | [km s] | |
2 | 262.6165032 | -34.6955865 | 0.065 | DCN, Single | ||||
3 | 262.6184156 | -34.6965240 | 0.040 | DCN, Single | ||||
4 | 262.6103159 | -34.6932659 | 0.155 | DCN, Single | ||||
5 | 262.6101515 | -34.6960014 | 0.087 | DCN, Single | ||||
6 | 262.6049155 | -34.6934384 | 0.126 | DCN, Single | ||||
7 | 262.6137738 | -34.6947298 | DCN & NH | |||||
8 | 262.6039531 | -34.6936374 | ||||||
9 | 262.6192359 | -34.6903650 | DCN & NH | |||||
11 | 262.6243189 | -34.6880780 | DCN & NH | |||||
12 | 262.6072148 | -34.6969795 | 0.019 | DCN, Single | ||||
13 | 262.6078228 | -34.6996836 | ||||||
14 | 262.6147937 | -34.6946762 | 0.090 | DCN, Single | ||||
15 | 262.6107433 | -34.6964412 | 0.087 | DCN, Complex | ||||
16 | 262.6215941 | -34.6989408 | ||||||
17 | 262.5954514 | -34.6916168 | ||||||
18 | 262.5927434 | -34.7052494 | ||||||
19 | 262.6064012 | -34.7019756 | ||||||
20 | 262.6111096 | -34.6932787 | ||||||
21 | 262.6131910 | -34.6939495 | 0.063 | DCN, Single | ||||
22 | 262.6118441 | -34.6946150 | DCN & NH | |||||
23 | 262.6028175 | -34.6925438 | ||||||
24 | 262.6198349 | -34.6960383 | 0.089 | DCN, Single | ||||
25 | 262.6155222 | -34.6952591 | DCN & NH | |||||
26 | 262.6143434 | -34.6917027 | ||||||
27 | 262.6000802 | -34.6910324 | ||||||
28 | 262.6253977 | -34.6999713 | ||||||
29 | 262.6133074 | -34.6919187 | ||||||
30 | 262.6114686 | -34.6962602 | 0.074 | DCN, Single | ||||
31 | 262.6096651 | -34.6925680 | ||||||
32 | 262.6094126 | -34.6910985 | ||||||
33 | 262.6287106 | -34.6862068 | ||||||
34 | 262.5982914 | -34.6919006 | 0.080 | DCN, Single | ||||
35 | 262.6142011 | -34.6940134 | DCN & NH | |||||
36 | 262.6202758 | -34.7001995 | ||||||
37 | 262.6010398 | -34.6950114 | 0.053 | DCN, Single | ||||
38 | 262.5971034 | -34.6920396 | ||||||
39 | 262.6054437 | -34.6963773 | DCN & NH | |||||
40 | 262.5917454 | -34.6897316 | ||||||
41 | 262.6095777 | -34.6983259 | ||||||
42 | 262.5975992 | -34.6876666 | ||||||
43 | 262.6035166 | -34.6966807 | 0.062 | DCN, Single | ||||
44 | 262.6030115 | -34.6956424 | 0.091 | DCN, Single | ||||
45 | 262.6143008 | -34.6909376 | ||||||
46 | 262.6187648 | -34.6912377 | DCN & NH | |||||
47 | 262.6178453 | -34.6919943 |
[…]
V-shape ID | (H) | Upper / lower VG | Mean VG | Upper / lower | (H) | |||
[ ] | [ ] | [M] | [km s pc] | [km s pc] | [kyr] | [kyr] | [10 M yr] | |
A | 353.3981 | -0.3506 | 20.85 / 17.04 | 46.89 / 57.39 | ||||
B | 353.4127 | -0.3632 | 25.34 / 17.22 | 38.59 / 56.77 | ||||
C | 353.4135 | -0.3657 | 17.69 / 13.26 | 55.28 / 73.75 | ||||
D | 353.4133 | -0.3727 | 12.85 / 21.22 | 76.07 / 46.07 | ||||
E | 353.4096 | -0.3521 | 12.47 / 3.63 | 78.38 / 269.57 | ||||
F | 353.4128 | -0.3604 | 15.75 / 15.61 | 62.09 / 62.64 | ||||
G | 353.4110 | -0.3630 | 22.59 / 24.46 | 43.29 / 39.97 | ||||
H | 353.4140 | -0.3627 | 21.28 / 39.68 | 45.94 / 24.64 | ||||
I | 353.4091 | -0.3595 | 22.34 / 11.22 | 43.77 / 87.17 |
Appendix F G353 power law density profile
Here we provide the derivation of the density profile used for the gravitationally collapsing sphere. We assume a power law density profile defined as:
where provides a good fit to the edges of the PV distribution seen in Fig. 8.
To determine the value of , we integrate this expression in a sphere (Eq. F.2), with pc. Based of different testings we set the total mass of the sphere to 150 M. We define which corresponds to the pixel size of the NH data at a distance of 2 kpc.
where ) = 150 M, pc, and = 5.65, from Eq. F.4, we obtain:

### v2

@@
-ALMA-IMF XV: NH kinematic analysis on the intermediate protocluster G353.41 Key Words.: stars: formation – ISM: clouds – ISM: kinematics and dynamics – ISM: molecules1 Introduction
+ALMA-IMF XIII: N2H+ kinematic analysis on the intermediate protocluster G353.41
+ALMA-IMF XIII: N2H+ kinematic analysis on the intermediate protocluster G353.41
 The ALMA-IMF Large Program provides multi-tracer observations of 15 Galactic massive protoclusters at matched sensitivity and spatial resolution.
-We focus on the dense gas kinematics of the G353.41 protocluster traced by NH (10), with a critical density of cm, and spatial resolution 0.02 pc.
-G353.41, at a distance of 2 kpc, is embedded in a larger scale (pc) filament and has a mass of M within pc.
-We extract the NH (10) isolated line component and we decompose it by fitting up to 3 Gaussian velocity components.
+We focus on the dense gas kinematics of the G353.41 protocluster traced by N2H+ (10), with an spatial resolution 0.02 pc.
+G353.41, at a distance of 2 kpc, is embedded in a larger scale (pc) filament and has a mass of M⊙ within pc2.
+We extract the N2H+ (10) isolated line component and we decompose it by fitting up to 3 Gaussian velocity components.
 This allows us to identify velocity structures that are either muddled or impossible to identify in the traditional position-velocity diagram.
 We identify multiple velocity gradients on large ( 1 pc) and small scales (0.2 pc).
-We find good agreement between the NH velocities and the previously reported DCN core velocities, suggesting that cores are kinematically coupled to the dense gas in which they form.
-We measure 9 converging “V-shaped” velocity gradients (km s pc) that are well-resolved (sizes pc), located in filaments, and are sometimes associated with cores near their point of convergence.
-The average timescales associated with the V-shapes are kyr, or about twice the free-fall time of cores in the same area (kyr) but substantially shorter than protostar lifetime estimates (0.5 Myr).
+We find good agreement between the N2H+ velocities and the previously reported DCN core velocities, suggesting that cores are kinematically coupled to the dense gas in which they form.
+We measure 9 converging “V-shaped” velocity gradients (VGs) (km s-1 pc-1) that are well-resolved (sizes pc), located in filaments, which are sometimes associated with cores near their point of convergence.
 We interpret these V-shapes as inflowing gas feeding the regions near cores (the immediate sites of star formation).
-We derive mass accretion rates in the range of M yr.
+We estimate the timescales associated with V-shapes as VG-1, and we interpret them as inflow timescales.
+The average inflow timescale is kyr, or about twice the free-fall time of cores in the same area (kyr) but substantially shorter than protostar lifetime estimates (0.5 Myr).
+We derive mass accretion rates in the range of M⊙ yr-1.
 This feeding might lead to further filament collapse and formation of new cores.
@@
 Thus these data are consistent with a comparatively slow global protocluster contraction under gravity, and faster core formation within, suggesting the formation of multiple generations of stars over the protocluster lifetime.
-ALMA-IMF XV: NH kinematic analysis on the intermediate protocluster G353.41
 Key Words.:
@@
 In this paper, we focus on the G353.41 protocluster (see Fig.
-1), and in particular, on the dense gas kinematics observationally accessible from the protocluster scale (2.9 pc) to the core scale.
-We trace this dense and cold gas using the NH (1-0) line observed with ALMA.
-Here NH, given its comparatively high critical density (see below), accesses the inner dense gas ”skeleton” of the protocluster structure, free from confusion induced by lower density gas.
+1), and in particular, on the dense gas kinematics observationally accessible from the protocluster scale (2.9 pc2) to the core scale.
+We trace this dense and cold gas using the N2H+ (1-0) line observed with ALMA.
+Given N2H+ is detected at high column densities (N(H2) cm-2; Tafalla et al., 2021), we gain access to the inner dense gas ”skeleton” of the protocluster structure, free from confusion induced by lower density gas.
 Meanwhile, ALMA permits us to obtain the resolution needed to trace structures down to the core scales where individual or small numbers of stars may be forming.
-The ALMA-IMF Large Program111Proposal ID 2017.1.01355.L, PIs: Motte, Ginsburg, Louvet, Sanhueza (LP) maps 15 dense, nearby ( kpc), and massive ( 10 M) Milky Way protoclusters down to 2 kau scales (Motte et al., 2022), at matched spatial resolution.
+The ALMA-IMF Large Program111Proposal ID 2017.1.01355.L, PIs: Motte, Ginsburg, Louvet, Sanhueza (LP) maps 15 dense, nearby ( kpc), and massive ( 103 M⊙) Milky Way protoclusters down to 2 kau scales (Motte et al., 2022), at matched spatial resolution.
 ALMA-IMF provides a large protocluster sample in order to test the universality of the stellar initial mass function (IMF) (Bastian et al., 2010; Offner et al., 2014).
@@
 This approach offers a thorough characterization of the processes taking place in these regions.
-Field size | Pixel scale | Beam size | BPA | RMS | Channel width | RMS velocity range | V |
-[°] | [K] | [ km s ] | [ km s ] | [ km s ] | |||
+Field size | Pixel scale | Beam size | BPA | aRMS | Channel width | RMS velocity range | bVLSR |
+[°] | [K] | [ km s-1 ] | [ km s-1 ] | [ km s-1 ] | |||
 176″172″ | 0.72″ | 1.96″2.29″ | 80.19 | 0.37 | 0.23 | [43 ; 32], [0 ; +7] | 17 |
@@
 In Fig.
-1 we indicate the ALMA-IMF NH (10) coverage of G353 (centered at , (J2000) = 17:30:26.28,34:41:49.7) and its parent filament (dark lane traced by ATLASGAL 870 m emission; Schuller et al., 2009) with light green and gray contours respectively.
+1 we indicate the ALMA-IMF N2H+ (10) coverage of G353 (centered at , (J2000) = 17:30:26.28,34:41:49.7) and its parent filament (dark lane traced by ATLASGAL 870 m emission; Schuller et al., 2009) with light green and gray contours respectively.
 Motte et al.
-(2022) classify this protocluster as being at an intermediate evolutionary state, located at 2 kpc, and hosting a total mass of 2.5 M.
+(2022) classify this protocluster as being at an intermediate evolutionary state, located at 2 kpc, and hosting a total mass of 2.5 M⊙.
 They describe G353 as isolated, without obvious interaction with massive nearby stellar clusters.
-Using moment maps derived from the NH (10) 12 m dataset they suggest the presence of multiple velocity components indicating a complex velocity field.
+Using moment maps derived from the N2H+ (10) 12 m dataset they suggest the presence of multiple velocity components indicating a complex velocity field.
 They propose that G353 is composed of filaments interacting at the central hub.
@@
 (2024), this region is an outlier in the ALMA-IMF hot core sample.
-Only one weak, low-mass ( 2 M) compact methyl formate source is detected and it lacks strong emission from complex organic molecules.
+Only one weak, low-mass ( 2 M⊙) compact methyl formate source is detected and it lacks strong emission from complex organic molecules.
 They state that this protocluster is in a chemically poor stage, where further characterization of this region is required.
-The NH (10) transition ( = 93.173809 GHz), given its high critical density, (Ungerechts et al., 1997), allows us to access the dense gas kinematics present in the innermost parts of star forming regions (Caselli et al., 2002a; Bergin et al., 2002; Tafalla et al., 2004; Lippok et al., 2013; Storm et al., 2014; Hacar et al., 2018; Chen et al., 2019; González Lobos & Stutz, 2019; Álvarez-Gutiérrez et al., 2021).
+The N2H+ (10) transition ( = 93.173809 GHz), given its high critical density, (Ungerechts et al., 1997), allows us to access the dense gas kinematics present in the innermost parts of star forming regions (Caselli et al., 2002a; Bergin et al., 2002; Tafalla et al., 2004; Lippok et al., 2013; Storm et al., 2014; Hacar et al., 2018; Chen et al., 2019; González Lobos & Stutz, 2019; Álvarez-Gutiérrez et al., 2021).
 The J = 10 transition presents seven hyperfine components (Cazzoli et al., 1985; Caselli et al., 1995, 2002a).
@@
 Such simplification is convenient to study the complex velocity fields found at the center of filaments.
-These regions present the densest environments for star formation, usually presenting multiple, blended velocity components, where the velocity distributions exhibit twists, turns, spirals, and wave-like patterns (Csengeri et al., 2011; Fernández-López et al., 2014; Stutz & Gould, 2016; Liu et al., 2019; González Lobos & Stutz, 2019; Álvarez-Gutiérrez et al., 2021; Sanhueza et al., 2021; Redaelli et al., 2022; Olguin et al., 2023).
+These regions present the densest environments for star formation, usually presenting multiple, blended velocity components, where the velocity distributions exhibit twists, turns, spirals, and wave-like patterns (Csengeri et al., 2011; Fernández-López et al., 2014; Stutz & Gould, 2016; Liu et al., 2019; González Lobos & Stutz, 2019; Henshaw et al., 2020; Álvarez-Gutiérrez et al., 2021; Sanhueza et al., 2021; Redaelli et al., 2022; Olguin et al., 2023).
 Recent techniques, such as the intensity-weighted position-velocity (PV) diagrams (González Lobos & Stutz, 2019; Álvarez-Gutiérrez et al., 2021), allow us to characterize processes such as infall, outflow, or rotation present in these environments, where high spatial and spectral resolution studies open a window into the small scale gas kinematics of star forming regions.
-In addition to the PV diagrams, we can create Position-Position-Velocity (PPV) diagrams, in order to identify coherent structures that might be both spatially and kinematically associated (Chen et al., 2019; Henshaw et al., 2019; Sanhueza et al., 2021; Redaelli et al., 2022).
-In this paper we investigate the NH dense gas kinematics of G353 from large (protocluster) to small (cores) scales.
+In addition to the PV diagrams, we can create Position-Position-Velocity (PPV) diagrams, in order to identify coherent structures that might be both spatially and kinematically associated (Chen et al., 2019; Henshaw et al., 2019, 2020; Sanhueza et al., 2021; Redaelli et al., 2022).
+In this paper we investigate the N2H+ dense gas kinematics of G353 from large (protocluster) to small (cores) scales.
 In § 2 we present the data.
-In § 3 we introduce our NH isolated extraction procedure.
-In § 4 we model and decompose the multiple velocity components found in the NH isolated component spectra.
+In § 3 we introduce our N2H+ isolated extraction procedure.
+In § 4 we model and decompose the multiple velocity components found in the N2H+ isolated component spectra.
 In § 5 we show our gas kinematic analysis, from protocluster to core scales.
 In § 6 we show that G353 might be under gravitational collapse at small and large scales.
-In § 7 we estimate mass accretion rates for multiple velocity gradients characterized in our NH data.
+In § 7 we estimate mass accretion rates for multiple velocity gradients characterized in our N2H+ data.
 We discuss our results in § 8, and we present our summary and conclusions in § 9.
@@
 2.1 ALMA-IMF data
-We make use of the NH (10) 12 m, 7 m, and Total Power observations described in Motte et al.
+We make use of the N2H+ (10) 12 m, 7 m, and Total Power observations described in Motte et al.
 (2022) for our analysis, providing robust uv plane coverage.
-We image the combination of the NH 7 m and 12 m (from now on called “7m+12m”) measurement set of G353, using the publicly available imaging scripts from the ALMA-IMF github repository333https://github.com/ALMA-IMF/reduction.
+We image the combination of the N2H+ 7 m and 12 m (from now on called “7m+12m”) measurement set of G353, using the publicly available imaging scripts from the ALMA-IMF github repository333https://github.com/ALMA-IMF/reduction.
 These data are corrected by the primary beam response pattern.
-Due to the missing large-scale emission, we find that near the V of the protocluster (17 km s; Wienen et al., 2015; Motte et al., 2022) some subregions in the 7m+12m cube present deep negative artifacts (“negative bowls”).
-To cover all possible uv scales, we combine the NH 7m+12m continuum-subtracted cube with the Total Power observations from the ALMA-IMF LP.
+Due to the missing large-scale emission, we find that near the VLSR of the protocluster (17 km s-1; Wienen et al., 2015; Motte et al., 2022) some subregions in the 7m+12m cube present deep negative artifacts (“negative bowls”).
+To cover all possible uv scales, we combine the N2H+ 7m+12m continuum-subtracted cube with the Total Power observations from the ALMA-IMF LP.
 We use the feather444https://casa.nrao.edu/docs/taskref/feather-task.html task from CASA 5.6.0.
@@
 To estimate and subtract the continuum emission present in the 7m+12m cube, we use the imcontsub555https://casa.nrao.edu/docs/taskref/imcontsub-task.html CASA task.
-We select the emission-free channels between 43 km s and 33 km s, and set the polynomial degree of the continuum fit (fitorder) to 0.
+We select the emission-free channels between 43 km s-1 and 33 km s-1, and set the polynomial degree of the continuum fit (fitorder) to 0.
 We list relevant final image parameters in Table 1, such as the field size, pixel scale, beam size, root-mean-square noise (RMS), and channel width.
 2.2 Core properties from published catalogues
-We use the 1.3 mm continuum derived cores catalogue666Available at www.almaimf.com, extracted using the getsf algorithm on the 1.3 mm continuum maps, smoothed at a common resolution of 2700 au (45 sources, Louvet et al.
-submitted).
-We also use the DCN core velocities (15 sources, Cunningham et al., 2023) and the SiO outflow catalogue (16 sources, Towner et al., 2024) in order to look for correlation between the NH gas kinematics and cores/outflows position and properties.
+We use the cores catalogue666Available at www.almaimf.com from Louvet et al.
+(submitted).
+These cores where identified using the getsf algorithm, specialized in source extraction on regions with complex filamentary structures (Men’shchikov, 2021).
+This procedure was done using the 1.3 mm continuum maps, smoothed at a common resolution of 2700 au, obtaining a total of 45 sources for G353.
+We also use the DCN core velocities (15 sources, Cunningham et al., 2023) and the SiO outflow catalogue (16 sources, Towner et al., 2024) in order to look for correlation between the N2H+ gas kinematics and cores/outflows position and properties.
 It is worth mentioning that, within a radius of from the center of G353 (Motte et al., 2022), we find 60% of the 1.3 mm cores (27 sources), and of the cores with DCN velocities and SiO outflows (11 sources from each catalogue).
@@
 The presence of these sources might imply a complex velocity field in this region, given that cores and outflows disturb the kinematics of the surrounding gas.
-3 NH isolated component extraction
-The NH (10) transition is characterized by its hyperfine emission composed by seven components (Caselli et al., 1995, see their Fig.
+3 N2H+ isolated component extraction
+The N2H+ (10) transition is characterized by its hyperfine emission composed by seven components (Caselli et al., 1995, see their Fig.
 1).
-We present an ideal example of NH emission in Fig.
+We present an ideal example of N2H+ emission in Fig.
 3, panel d.
-In this work we refer to the triplet of hyperfine components that present the highest intensities as the main NH components, located at the center of the line emission at GHz.
-We refer to the most blueshifted hyperfine component as the isolated component, at GHz, shifted by km s relative to the main NH component (see Table 1 from Cazzoli et al., 1985).
+In this work we refer to the triplet of hyperfine components that present the highest intensities as the main N2H+ components, located at the center of the line emission at GHz.
+We refer to the most blueshifted hyperfine component as the isolated component, at GHz, shifted by km s-1 relative to the main N2H+ component (see Table 1 from Cazzoli et al., 1985).
 We developed an algorithm to extract only the isolated hyperfine component from every pixel in the feathered datacube.
 This is in order to reduce the complexity of our data, given that it may contain multiple velocity components in addition to the hyperfine line emission.
-Considering that the NH emission moves in velocity across the protocluster, our approach is to find the velocity where the emission of the isolated component ends and remove the rest of the line emission.
-We also preserve the emission-free channels, at low ( km s to km s) and high ( km s to km s) velocities, to improve future RMS estimations if needed.
-Note that in the procedures described below, we use findpeaks777https://docs.scipy.org/doc/scipy/reference/generated/scipy.signal.find_peaks.html to detect peaks and valleys in different spectra.
-Our extraction approach is separated into two procedures, for low and for high signal-to-noise (SNR) data.
-In order to determine which data have low or high SNR, we obtain the mean spectrum over all the spatial pixels of the cube, which serves as a guide to determine the velocity at the “mean dip” (V = -22 km s, dashed gray line in Fig.
+Considering that the N2H+ emission moves in velocity across the protocluster, our approach is to find the velocity where the emission of the isolated component ends and remove the rest of the line emission.
+We also preserve the emission-free channels, at low ( km s-1 to km s-1) and high ( km s-1 to km s-1) velocities, to improve future RMS estimations if needed.
+Note that in the procedures described below, we use findpeaks777https://docs.scipy.org/doc/scipy/reference/generated/scipy.signal.find_peaks.html to detect peaks and valleys in the different spectra.
+Our extraction approach is separated into two procedures, for low and for high signal-to-noise (SNR) data (see Fig.
+2, and text below).
+In order to determine which data have low or high SNR, we obtain the mean spectrum over all the spatial pixels of the cube, which serves as a guide to determine the velocity at the “mean dip” (V = -22 km s-1, dashed gray line in Fig.
 3 panel “a”).
-This velocity represents the mean location of the intensity valley between the isolated and the main components of the NH emission.
-We define V = 3.2 km s as the difference between V and the velocity at the peak of the mean isolated component V (dashed black line in Fig.
+This velocity represents the mean location of the intensity valley between the isolated and the main components of the N2H+ emission.
+We define Vmean = 3.2 km s-1 as the difference between V and the velocity at the peak of the mean isolated component V (dashed black line in Fig.
 3 panel “a”), used in our velocity guesses for the high SNR extraction procedure (see below).
-To create a SNR map of the isolated component, we first measure the RMS noise in emission-free channels ( km s to km s), and the peak intensity in the channels range where the mean isolated component is located ( km s to V).
+To create a SNR map of the isolated component, we first measure the RMS noise in emission-free channels ( km s-1 to km s-1), and the peak intensity in the channels range where the mean isolated component is located ( km s-1 to V).
 This approach allows us to exclude the emission of the main line components.
@@
 Then, we use binarypropagation on the cleaned SNR map, using the original SNR map mask, to restore only the protocluster edges.
-To test our cleaning approach we compute the total integrated intensity using the Python package SpectralCube101010https://github.com/radio-astro-tools/spectral-cube in the range of 31.5 km s to V using the original and cleaned SNR mask.
+To test our cleaning approach we compute the total integrated intensity using the Python package SpectralCube101010https://github.com/radio-astro-tools/spectral-cube in the range of 31.5 km s-1 to V using the original and cleaned SNR mask.
 We estimate that the removed spurious emission accounts for 2% of the total integrated intensity for data with SNR.
 In Fig.
-2 we show the NH isolated component SNR map, where at SNR values 5 we capture the cloud emission while excluding noise (white contour).
+2 we show the N2H+ isolated component SNR map, where at SNR values 5 we capture the cloud emission while excluding noise (white contour).
 We set our isolated component SNR threshold to 5, in order to use one of the two extraction procedures (see below).
 In this section we refer to high (low) SNR spectrum if its isolated component ().
-For low SNR spectra, we extract all the channels in the velocity range from km s up until V (panel “b” in Fig.
+For low SNR spectra, we extract all the channels in the velocity range from km s-1 up until V (panel “b” in Fig.
 3).
+We take this approach given that for low SNR data we can not identify peaks in the N2H+ emission in a reliable manner.
 For high SNR spectra the extraction procedure consists of creating different velocity guesses that represent the location of the intensity valley, similar to the definition of V (Fig.
@@
 This approach is described in detail here:
-First, we implement a rolling average along each spectra.
+First, we implement a rolling average along each spectrum.
 This is in order to smooth over intensity bumps that might result in false positives for the detection of peaks and valleys.
 For this procedure, we average considering two channels before and after each velocity.
-After smoothing, for each spectra we identify the isolated component peak using find_peaks.
+After smoothing, for each spectrum we identify the isolated component peak using find_peaks.
 We call the velocity associated to this peak V.
 In the case of multiple velocity components it represents the most blueshifted one.
-We find the intensity valley between the isolated component and the NH main line emission by inverting the spectra and finding the first peak which is the inverted intensity valley.
+We find the intensity valley between the isolated component and the N2H+ main line emission by inverting the spectrum and finding the first peak which is the inverted intensity valley.
 We define the associated velocity to this intensity valley as V.
 We create three velocity guesses based on the properties of each spectrum in our cube (see points below).
-These are the 1 guess: VV.
-2 guess: V, In the case of multiple isolated components this guess might incorrectly capture the intensity valley after the first isolated component.
+These are the 1st guess: VVmean.
+2nd guess: V, In the case of multiple isolated components this guess might incorrectly capture the intensity valley after the first isolated component.
 In that case, the other guesses are needed for a reliable isolated component extraction.
-3 guess: VV to provide a velocity cut further away from the V.
+3rd guess: VVmean to provide a velocity cut further away from the V.
 This guess is mainly useful in the case where multiple isolated components cover a velocity range larger than the one probed by the other two guesses.
@@
 We do not encounter divergences in this normalization given these parameters are not exactly zero.
-(1) where the “norm” subscript indicates that the parameter list is divided by its minimum value.
+(1) where the “norm” subscript indicates that the parameter list is normalized by dividing it by its minimum value.
 By visual inspection we consider that we obtain good extraction results when the weight is mostly dependent on and in a minor part on .
@@
 We choose the guess with the weight closest to unity.
-Similarly as for SNR 5, we extract the spectra from km s up until the velocity of the chosen guess, and preserving the emission-free channels from 0.7 km s to 6.7 km s.
-Various examples of NH spectra and isolated hyperfine component extraction are shown in Fig.
-3, where we can see spectra containing one (panel “b”), two (panel “c”), and three (panel “d”) velocity components, all well extracted by our procedure.
+Similarly as for SNR 5, we extract the spectra from km s-1 up until the velocity of the chosen guess, and preserving the emission-free channels from 0.7 km s-1 to 6.7 km s-1.
+Various examples of N2H+ spectra and isolated hyperfine component extraction are shown in Fig.
+3, where we can see spectra containing one (panel “c”), two (panel “e”), and three (panel “f”) velocity components, all well extracted by our procedure.
 In Stutz et al.
-(in prep) this approach is generalized to all ALMA-IMF regions for NH, providing reliable results.
+(in prep) this approach is generalized to all ALMA-IMF regions for N2H+, providing reliable results.
+3.1 Filamentary identification
+We use the FilFinder Python package (Koch & Rosolowsky, 2015) in order to detect the most prominent filaments in this region (green lines in Fig.
+4).
+The procedure, including the parameters we used for the filamentary identification, is presented in Appendix A.
 In Fig.
-4 we show the moment 0 map estimated over the extracted NH isolated velocity components.
-We use the FilFinder Python package (Koch & Rosolowsky, 2015) in order to identify the most prominent filaments in this region (black lines in Fig.
-4).
+4 we indicate the detected filaments with green lines, on top of the moment zero map of the multiple N2H+ isolated components.
+We see that G353 is a hub-filament system (HFS), composed by three main filaments converging at its center.
+The HFSs are a characteristic feature of early stages of star formation, where gas flows through the filaments towards the central hub triggering star formation (Myers, 2009; Galván-Madrid et al., 2010; Busquet et al., 2013; Galván-Madrid et al., 2013; Peretto et al., 2014; Kumar et al., 2022; Zhou et al., 2022).
 We see that in the plane of the sky (POS) most of the 1.3 mm cores (red ellipses) are located on top of the filaments.
 This spatial agreement between filaments and protostelar cores is consistent with filamentary fragmentation (André et al., 2010; Busquet et al., 2013; Stutz & Kainulainen, 2015; Kuznetsova et al., 2015, 2018).
-The procedure, including the parameters we used for the filamentary identification, is presented in Appendix A.
-4 NH isolated component velocity decomposition
+4 N2H+ isolated component velocity decomposition
 In Fig.
 3 we see that clear multiple isolated velocity components are present in our dataset.
-To characterize the complex dense-gas kinematics traced by NH we follow the method in Álvarez-Gutiérrez et al.
+To characterize the complex dense-gas kinematics traced by N2H+ we follow the method in Álvarez-Gutiérrez et al.
 (2021), and we use the spectroscopic toolkit PySpecKit (Ginsburg & Mirocha, 2011; Ginsburg et al., 2022) to model and decompose the isolated component emission.
-PySpecKit adjusts a fixed number of components set by the user, based on visual inspection of the data we impose three velocity components to every spectra and then remove false positives (see below).
+PySpecKit adjusts a fixed number of components set by the user, based on visual inspection of the data we impose three velocity components to every spectrum and then remove false positives (see below).
 Given the kinematic complexity of the data and cursory inspection of the spectra, a simpler analysis with only two components contradicts the data.
@@
 To improve the convergence of PySpecKit, we create a set of ranges for the parameters that define each of the three Gaussian velocity components, namely the peak intensity, central velocity, and velocity dispersion.
-After testing different parameter ranges, we set the intensity range between 1.76 K (4 times the mean RMS) and 30 K, the velocity centroid from 30 km s to 20 km s, and the velocity dispersion from 0.22 km s to 1 km s.
+After testing different parameter ranges, we set the intensity range between 1.76 K (4 times the mean RMS) and 30 K, the velocity centroid from 30 km s-1 to 20 km s-1, and the velocity dispersion from 0.22 km s-1 to 1 km s-1.
 From the results using the ranges defined above, we notice that some modeled components do not fit any emission.
@@
 We present the Gaussian fits of the high SNR spectra from Fig.
-3 in Fig.
-B.1.
+3 in Appendix B.
 In Fig.
 5 we show the spatial distribution of the multiple Gaussian velocity components.
-In gray we indicate the main filamentary structure in the region (see § 3).
+In gray we indicate the main filamentary structure in the region (see § 3.1).
 The first and second velocity components, in blue and green respectively, present most of the high intensity emission and they also spatially dominate over the third, most red-shifted component.
@@
 In Fig.
-6 we present the number of Gaussian velocity components for each spectra, where we highlight that:
-Most of the NH data presents emission characterized by two velocity components.
+6 we present the number of Gaussian velocity components for each spectrum, where we highlight that:
+Most of the N2H+ data presents emission characterized by two velocity components.
 Most of the spectra described by three velocity components are located in the innermost parts of the region.
 Most of the cores (black ellipses; Louvet et al.
-submitted) are located in regions with spectra presenting two to three Gaussian velocity components, indicating kinematic complexity even at 4 kau (NH spatial resolution).
+submitted) are located in regions with spectra presenting two to three Gaussian velocity components, indicating kinematic complexity even at 4 kau (N2H+ spatial resolution).
 Single velocity component spectra are located preferentially in the outskirt of the protocluster.
@@
 7 we show the histogram of the fitted velocity centroid of each Gaussian velocity component.
-The peaks of these distributions are located at 27, 24.7, and 23.3 km s respectively, well-separated in velocity.
+The peaks of these distributions are located at 27, 24.7, and 23.3 km s-1 respectively, well-separated in velocity.
 From hereafter we refer to these distribution as blue, green, and red respectively.
 Most of the velocity components appear to be associated with the blue and green distributions.
-For consistency with the different tracers used in further analysis, we shift the isolated component velocities by +8 km s, to the reference frame of the main line components of NH (Cazzoli et al., 1985).
-4.1 DCN & NH derived core velocities
+For consistency with the different tracers used in further analysis, we shift the isolated component velocities by +8 km s-1, to the reference frame of the main line components of N2H+ (Cazzoli et al., 1985).
+4.1 DCN & N2H+ derived core velocities
 In this section our goal is to increase the sample of core velocities from the already published DCN catalog, aiming to explore all the potential in these types of dataset.
-Given the relatively high of DCN (32) (cm) compared to NH (10) (), DCN (32) is known to coincide well with continuum peaks associated to cores (Liu et al., 2015; Cunningham et al., 2016; Minh et al., 2018), while NH is characterized by tracing the dense gas at the innermost parts of star forming regions (Fernández-López et al., 2014; Hacar et al., 2018; González Lobos & Stutz, 2019).
+Given the relatively high of DCN (32) (cm-3) compared to N2H+ (10) (), DCN (32) is known to coincide well with continuum peaks associated to cores (Liu et al., 2015; Cunningham et al., 2016; Minh et al., 2018), while N2H+ is characterized by tracing the dense gas at the innermost parts of star forming regions (Fernández-López et al., 2014; Hacar et al., 2018; González Lobos & Stutz, 2019).
 In Cunningham et al.
@@
 We use the ALMA-IMF DCN 12 m data from Cunningham et al.
-(2023), which presents a velocity resolution of km s.
-For each DCN velocity core described by a single component (Cunningham et al., 2023) we compare the emission of the DCN and modeled NH isolated spectra.
-We find an average velocity offset between the DCN peak and the closest NH isolated component peak of km s, less than two DCN channel widths.
+(2023), which presents a velocity resolution of km s-1.
+For each DCN core velocity described by a single component (Cunningham et al., 2023) we compare the emission of the DCN and modeled N2H+ isolated spectra.
+We find an average velocity offset between the DCN peak and the closest N2H+ isolated component peak of km s-1, less than two DCN channel widths.
 We use this approach for the remaining 30 cores, in order to determine their DCN velocities.
-Here we estimate the RMS of the DCN data in emission-free channels in the range of 42 km s to 25 km s and we obtain the SNR map by dividing the peak intensity by the RMS.
+Here we estimate the RMS of the DCN data in emission-free channels in the range of 42 km s-1 to 25 km s-1 and we obtain the SNR map by dividing the peak intensity by the RMS.
 For the procedure below we only use DCN spectra with SNR 3.
-Next, we extract the average DCN and modeled NH isolated component spectra of these 30 cores.
-We identify the NH isolated velocity component closest to the DCN peak within three DCN channel widths.
-We find that 11 out of these 30 cores present DCN with SNR close to one NH velocity component.
+Next, we extract the average DCN and modeled N2H+ isolated component spectrum of these 30 cores.
+We identify the N2H+ isolated velocity component closest to the DCN peak within three DCN channel widths.
+We find that 11 out of these 30 cores present DCN with SNR close to one N2H+ velocity component.
 Here, we define the velocity of these cores as the velocity where the DCN emission peaks.
-On average, these cores have a velocity offset between these two tracers less than 0.8 km s ( DCN channels), similar to the results obtained for the 15 cores with DCN velocities from Cunningham et al.
-(2023), and they present an average velocity offset of 0.38 km s.
-Throughout this paper we refer to these cores as “DCN & NH cores” given they are derived from the comparison of these two tracers.
-In Table C.1 we include these DCN & NH core velocities, complementing the DCN catalogue from Cunningham et al.
+On average, these cores have a velocity offset between these two tracers less than 0.8 km s-1 ( DCN channels), similar to the results obtained for the 15 cores with DCN velocities from Cunningham et al.
+(2023), and they present an average velocity offset of 0.38 km s-1.
+Throughout this paper we refer to these cores as “DCN & N2H+ cores” given they are derived from the comparison of these two tracers.
+In Appendix C we present two examples of the comparison of the DCN and N2H+ spectra in cores where we see clear agreement between these tracers (Fig.
+C.1).
+In Table C.1 we show these obtained core velocities from our comparison between DCN & N2H+ spectra, complementing the DCN catalogue from Cunningham et al.
 (2023).
-In Fig.
-C.1 we present two examples of DCN and NH spectra in cores where we see clear agreement between these tracers.
 5 Analysis of position-velocity diagrams
@@
 We create this diagram by taking the total intensity along the Galactic longitude, where indicates the distance in parsec relative to the center of G353, assuming a distance to the protocluster of of 2 kpc (Motte et al., 2022).
-We see general agreement between the DCN core velocities and the NH velocity distribution.
+We see general agreement between the DCN core velocities and the N2H+ velocity distribution.
 This suggests that most of the cores are still kinematically coupled to the dense gas in which they formed.
-As presented in § 4.1, the DCN and NH velocities match within 0.8 km s ( DCN channels).
+As presented in § 4.1, the DCN and N2H+ velocities match within 0.8 km s-1 ( DCN channels).
 […]
 Regarding the dense gas velocity distribution, in Fig.
-8 we see a velocity spread of km s in the sub-region between 0.3 pc to 0.1 pc.
+8 we see a velocity spread of km s-1 in the sub-region between 0.3 pc to 0.1 pc.
 Most of the intensity on this diagram is located at the upper part of this sub-region, at .
@@
 9 we show the spatial distribution of the fitted Gaussian velocity components (see § 4).
-The blue, green, and red color maps indicate the integrated intensity of the first, second, and third velocity components of the NH spectra respectively.
+The blue, green, and red color maps indicate the integrated intensity of the first, second, and third velocity components of the N2H+ spectra respectively.
 Note that the spatial overlap between any of these components is presented in Fig.
@@
 Using these parameters we create intensity-weighted PV diagrams along the and coordinates.
-We present these NH PV diagrams in the bottom left and top right panels of Fig.
+We present these N2H+ PV diagrams in the bottom left and top right panels of Fig.
 9.
 The key features on the position-position (PP) and on the top right PV diagram, are:
-The agreement between the DCN core velocities and the overall NH PV structures suggests that cores are still kinematically coupled to the dense gas in which they formed.
-We see at least nine clear and prominent V-shaped velocity gradients (see Fig.
-D.1), across all velocity components.
+The agreement between the DCN core velocities and the overall N2H+ PV structures suggests that cores are still kinematically coupled to the dense gas in which they formed.
+We see at least nine clear and prominent V-shaped velocity gradients (see Appendix D), across all velocity components.
 The orientation of these V-shape, pointing to the left/right (top right panel) or up/down (bottom left panel), follow no clear preference.
@@
 This technique recovers the large scale velocity spread present in Fig. 8 and highlights small scale structures.
-The most prominent V-shape is located at pc, 20.5 km s), between two 1.3 mm cores with DCN detections (see § 6).
+The most prominent V-shape is located at pc, 20.5 km s-1), between two 1.3 mm cores with DCN detections (see § 6).
 For a better visualization of the 3D structure of these V-shapes we provide an interactive 3D PPV diagram at: rodrigoalvarez.space/research/figures.
@@
 Given the visual linearity of the VGs composing the V-shape, we apply a linear fit to these distribution in order to characterize them.
-For these fits we consider data only above an integrated intensity threshold of 8 K km s and 3 K km s, for the upper and lower gradient respectively.
+For these fits we consider data only above an integrated intensity threshold of 8 K km s-1 and 3 K km s-1, for the upper and lower gradient respectively.
 We remove data not related to the velocity gradient, clustered in the ranges of , which lie just outside the filament hosting this V-shape on the POS.
 Additionally, we weight each point based on their integrated intensity to make our fits more robust.
-The slopes of the linear fits represent the VGs in km s pc.
+The slopes of the linear fits represent the VGs in km s-1 pc-1.
 These linear fits follow the VGs distribution and these are somewhat asymmetric, the upper gradient is slightly shallower than the bottom gradient.
@@
 These are related as VG = VG.
-These VGs present values between to km s pc (see Fig.
+These VGs present values between to km s-1 pc-1 (see Fig.
 10).
 Additionally, we estimate the center of this V-shape as the velocity-weighted mean position of the points composing this structure.
-With this approach the position of the points closest to the V-shape apex present more weight, obtaining the center of this V-shape at () = (353.4135, 0.3657).
+With this approach the position of the points closest to the V-shape apex present more weight, obtaining the center of this V-shape at () = (353.4135∘, 0.3657∘).
 This position is located between “core 2” and “core 3”, both of them having DCN velocity fits (Cunningham et al., 2023).
-These core present masses of 20.7 and 6.4 M respectively.
+These core present masses of 20.7 and 6.4 M⊙ respectively.
 We inspect the core catalog derived from the map at native resolution (Louvet et al.
@@
 In the left panel of Fig.
-11 we show the integrated intensity of the multiple modeled NH isolated components.
+11 we show the integrated intensity of the multiple modeled N2H+ isolated components.
 With colored boxes we show the areas where we create the different PV diagrams presented on the right panel.
@@
 We show that the overall structure in PV space is conserved at different angles, excluding the possibility of this velocity feature being the result of projection effects.
-For this V-shaped structure, we note that the NH velocity range is V 2 km s, while the whole NH isolated component dataset present a V 8 km s (Fig.
-9).
-From these VGs we derive timescales as = 1/VG, similar to the procedure for a rotating filament presented in Álvarez-Gutiérrez et al.
-(2021).
-We obtain timescales between kyr.
-These timescales are short compared to the 0.21 Myr free fall time (t) of the protocluster (Motte et al., 2022), and a few times larger than the t of nearby cores ( 20 kyr, within 0.1 pc of this V-shape).
-To determine the cores t, we use the 1.3 mm core masses from Louvet et al.
+For this V-shape we use its composing VGs to derive timescales as = 1/VG, similar to the procedure for a rotating filament presented in Álvarez-Gutiérrez et al.
+(2021), and we suggest these can be interpreted as inflow timescales.
+The values for this V-shape range between kyr.
+These timescales are short compared to the 0.21 Myr free fall time (tff) of the protocluster (Motte et al., 2022), and a few times larger than the tff of nearby cores ( 20 kyr, within 0.1 pc of this V-shape).
+To determine the cores tff, we use the 1.3 mm core masses from Louvet et al.
 (submitted).
-We characterized eight more NH V-shaped structure.
-These are presented in Appendix D.
+We characterized eight more N2H+ V-shaped structures.
+In Fig.
+13 we show the position of these V-shapes in the POS.
+Previous studies have detected velocity gradients along filaments, towards a converging point (Peretto et al., 2014; Pan et al., 2024; Rawat et al., 2024).
+In the case of G353 we see that instead of detecting a single V-shape at the intersection of its filaments, the hub appears to be fragmented into multiple, small-scale V-shaped VG.
+Only V-shapes “A”, “D”, and “E” are outside of the hub, with V-shape “D” located on top of filament “F3”.
+This indicates that the V-shaped structures are not exclusive to the central parts of HFS, but are also present in comparatively isolated regions.
 Note that within a beam size from the apex of V-shape “B” (see Fig.
-D.2), the continuum core “7” (M) is located.
+D.2), the continuum core “7” (M⊙) is located.
 In Henshaw et al.
@@
 E.1 (see Appendix E for more details).
-We note that there is almost no SiO emission nor outflow sources at the location of the NH V-shape (see Fig.
+We note that there is almost no SiO emission nor outflow sources at the location of the N2H+ V-shape (see Fig.
 12).
-The V within 10 from this main blue V-shape is 40 km s, while for the whole SiO dataset is 80 km s, showing a clear difference in the SiO V and the core velocities.
-Furthermore, the SiO V is times larger than the one of NH.
+The V within 10′′ from this main blue V-shape is 40 km s-1, while for the whole SiO dataset is 80 km s-1, showing a clear difference in the SiO V and the core velocities.
+Furthermore, the SiO V is times larger than the one of N2H+.
 This difference in traced velocities implies that these two molecules trace vastly different physical phenomena.
-We suggest that the small velocity range probed by NH indicates that the velocity gradients can be considered as infall signatures (see § 6).
+We suggest that the small velocity range probed by N2H+ indicates that the velocity gradients can be considered as infall signatures (see § 6).
 We discuss the possible morphology of the filaments hosting V-shapes in § 8.
@@
 10.
-For CO we measure the RMS in the emission-free velocity range of 145 to 286 km s.
+For CO we measure the RMS in the emission-free velocity range of 145 to 286 km s-1.
 We use only CO data with SNR3 for our analysis.
@@
 12).
-We find that the difference between the center of the V-shape and the barycenter of these cores is ″(au), well below the beam size of the NH data.
+We find that the difference between the center of the V-shape and the barycenter of these cores is ″(au), well below the beam size of the N2H+ data.
 A similar offset is also present in the intensity and velocity profiles along filaments from ATOMS data (Zhou et al., 2022, see their Fig.
 6).
-This small spatial offset might suggest that the gas flowing in the V-shape is produced by the gravitational pull towards the barycenter of cores 2 & 3, where the NH radial velocities peak.
+This small spatial offset might suggest that the gas flowing in the V-shape is produced by the gravitational pull towards the barycenter of cores 2 & 3, where the N2H+ radial velocities peak.
 This interpretation is similar to the one proposed in Zhou et al.
@@
 In Fig.
-13 we show the mean spectra of different tracers at the central position of the main blue V-shape.
-These spectra are measured over a circular region with diameter equal to the major axis of the NH beam (2.28 ″; ).
-This circular area results in a coverage of 1.14 times the NH beam, and 5.6 times the beams of the HCO, DCN, and HCO data.
-We see NH and HCO present double component spectra with asymmetric peaks.
-Between these peaks we detect DCN and HCO emission.
-The asymmetric spectra present in NH and HCO is consistent with the “blue asymmetry” spectral feature, usually interpreted as infall signature, suggesting that this region is under gravitational collapse (e.g.
+14 we show the mean spectrum of different tracers at the central position of the main blue V-shape.
+These spectra are measured over a circular region with diameter equal to the major axis of the N2H+ beam (2.28 ″; ).
+This circular area results in a coverage of 1.14 times the N2H+ beam, and 5.6 times the beams of the H2CO, DCN, and H213CO data.
+We see N2H+ and H2CO present double component spectrum with asymmetric peaks.
+Between these peaks we detect DCN and H213CO emission.
+The asymmetric spectrum present in N2H+ and H2CO is consistent with the “blue asymmetry” spectral feature, usually interpreted as infall signature, suggesting that this region is under gravitational collapse (e.g.
 Anglada et al., 1987; Mardones et al., 1997; Lee et al., 1999, 2001; Smith et al., 2012).
@@
 Below we compare this velocity spread with the velocity distribution produced by infall, where the gas velocities increase as the distance to the center of infall (“”) decreases:
-For this comparison we model a sphere with a total mass of 150 M, a radius of 0.5 pc, and a power law density profile described by:
+For this comparison we model a sphere with a total mass of 150 M⊙, a radius of 0.5 pc, and a power law density profile described by:
 we provide the derivation of in Appendix F.
@@
 We obtain the radial component of the infalling velocities as:
-where represents the horizontal coordinate in the POS, while represents the (non-observed) depth of the sphere.
+where represents the horizontal coordinate in the POS, while represents the (non-observed) depth of the sphere. The spatial coordinates for this model range from pc to pc.
 In Fig.
-14, we show the coverage of the PV distribution from our model with a solid white line.
+15, we show the coverage of the PV distribution from our model, described in Eqs.
+2 4, with a solid white line.
 We find good agreement between the PV distributions of our approach and the data.
@@
 At large scales we interpret the agreement between the PV diagrams of our model and the data as protocluster scale collapse due to gravitational contraction.
-It is worth noting that the mass of our model 5.5 times lower than the mass derived from the (H) map (Díaz-González et al., 2023).
+It is worth noting that the inferred mass from our model is 5.5 times lower than the one derived from the (H2) map (Díaz-González et al., 2023).
 We speculate that a model considering complex processes such as turbulence, radiative transfer, and magnetic fields might solve this mass discrepancy while still matching the observed PV distribution.
 7 Mass accretion rates in the V-shaped structure
-Based on the idea that the V-shapes are a result of gas flowing toward cores, in this section we provide estimates of their mass accretion rates () for NH and H.
-7.1 NH mass accretion rate
-To derive the associated to the main blue V-shape (Fig.
-10), we need to estimate its NH mass.
-For this procedure we use PySpecKit with the n2hp_vtau fitter, to fit the full NH line.
-The fitted parameters (see below) allow us to derive the (NH).
-The V-shape structure contains 77, 143, and 25 NH spectra with one, two, and three velocity components respectively.
-The bluest velocities in the three velocity component spectra accounts for 2% of the total number of velocities in this V-shape.
-For this reason, we model the full NH hyperfine line structure with one and two velocity components.
-In Table 2 we list the parameters and ranges used for this procedure.
-After obtaining the modeled NH cube, we remove modeled components where , where and represent the estimated opacity and its associated error, respectively.
-This criterion is to ensure that we use reliable fitted parameters to determine our (NH) values.
-For the fitting of two NH components we only analyze the most blueshifted component.
-The resulting opacities follow a log-normal distribution with a peak at .
-We derive the (NH) of the V-shape by using Eq. 5 (Caselli et al., 2002b):
-Where , T, and are the opacity, excitation temperature, and velocity dispersion respectively, obtained from the full line fitting.
-The Planck and Boltzmann constants are represented by and respectively, and are the frequency and wavelength of NH, is the Einstein coefficient of the NH (10) transition, and are the statistical weights of the lower and upper energy levels, is the partition function, and is the energy of the lower level.
-From the above procedure, inside the V-shaped structure, we get a total (NH) = 5.2410 cm and a total (NH) = 5.910 M.
-We use the average timescale of the VGs from § 5.3 (Fig.
-10), = 64.5 kyr, to determine the (NH) as:
-Note that the (NH) estimate (and (H) below) should be multiplied by , in order to account for the unknown inclination angle () of the protocluster/filaments relative to the POS.
-Excitation temperature [K] | (T) | 2.73 80 |
-Opacity | () | 0.01 40 |
-Centroid velocity [km s] | () | 25 15 |
-Velocity dispersion [km s] | () | 0.20 3 |
-7.2 H mass accretion rate
-To ensure that we estimate the V-shape (H) on the same area as in the NH hyperfine line fitting, we use the CASA task imregrid to obtain the continuum-derived (H) map from Díaz-González et al.
-(2023) at the resolution of the NH data.
-We determine that in this V-shape the total (H 1.03 cm.
-Here, we derive a (H) map using Eq.
-7:
-where from this (H) map we consider only the points that are part of the V-shape.
+Based on the idea that the V-shapes are a result of gas flowing toward cores, in this section we provide estimates of their mass accretion rates () for N2H+ and H2.
+7.1 H2 mass accretion rate
+To ensure that we estimate the V-shape (H2) on the same area as in the N2H+ hyperfine line fitting, we use the CASA task imregrid to obtain the continuum-derived (H2) map from Díaz-González et al.
+(2023) at the resolution of the N2H+ data.
+We determine that in this V-shape the total (H 1.17 cm-2 in an area of 0.013 pc2.
+Here, we derive a (H2) map using Eq.
+5:
+where from this (H2) map we consider only the points that are part of the V-shape.
 To determine the mass associated to flowing motions we subtract the core masses (from Louvet et al.
@@
 Note here that this mass map is an upper limit given that we do not apply a background correction.
-We obtain a total of (H 54 M.
+We obtain a total of (H 53 M⊙.
 Considering used in Eq.
-6, we derive the (H) as:
-We use the procedure described in this section to estimate the (H) of other eight V-shapes shown in Fig.
+8, we derive the (H2) as:
+We use the procedure described in this section to estimate the (H2) of other eight V-shapes shown in Fig.
 D.2.
 We include these values in Table D.1.
-The average (H) of these V-shapes is 3.4 10 M yr.
-We estimate the free-fall time of all 45 1.3 mm cores from Louvet et al.
-(submitted).
-These values present large scattering, ranging from M yr, with 28 of these cores presenting (H) M yr.
-For cores 2 & 3, the average (H) is 15.5 M yr, about twice the (H) of the main blue V-shape, located between these two cores.
-7.3 NH relative abundance
-For the main blue V-shape (Fig. D), we derive the NH relative abundance (NH), using the NH and H column densities obtained above, as:
-The NH) value obtained above is consistent, within the scatter, with other estimates in massive Galactic star formation regions in the ranges of (1.6 - 3.8) 10 (Caselli et al., 2002a; Henshaw et al., 2014, Sandoval-Garrido et al.
+The average (H2) of these V-shapes is 3.4 10-4 M⊙ yr-1.
+Note that V-shapes “H”, “C”, “F”, and “B” present the largest (H2) and they are located near or at the convergence point of the filaments (see Fig.
+13)
+For comparison, using the core masses from Louvet et al.
+(submitted) we estimate the free-fall time of all 45 1.3 mm cores and their mass accretion rates.
+These values present large scattering, ranging from M⊙ yr-1, with 28 of these cores presenting (H2) M⊙ yr-1.
+For cores 2 & 3, the average (H2) is 15.5 M⊙ yr-1, about twice the (H2) of the main blue V-shape, located between these two cores.
+7.2 N2H+ mass accretion rate and relative abundance
+To derive the associated to the main blue V-shape (Fig.
+10), we need to estimate its N2H+ mass.
+For this procedure we use PySpecKit with the n2hp_vtau fitter, to fit the full N2H+ line.
+The fitted parameters (see below) allow us to derive the (N2H+).
+The V-shape structure contains 77, 143, and 25 N2H+ spectra with one, two, and three velocity components respectively.
+The bluest velocities in the three velocity component spectra accounts for 2% of the total number of velocities in this V-shape.
+For this reason, we model the full N2H+ hyperfine line structure with one and two velocity components.
+In Table 2 we list the parameters and ranges used for this procedure.
+After obtaining the modeled N2H+ cube, we remove modeled components where , where and represent the estimated opacity and its associated error, respectively.
+This criterion is to ensure that we use reliable fitted parameters to determine our (N2H+) values.
+For the fitting of two N2H+ components we only analyze the most blueshifted component.
+The resulting opacities follow a log-normal distribution with a peak at .
+We derive the (N2H+) of the V-shape by using Eq. 7 (Caselli et al., 2002b):
+Where , Tex, and are the opacity, excitation temperature, and velocity dispersion respectively, obtained from the full line fitting.
+The Planck and Boltzmann constants are represented by and respectively, and are the frequency and wavelength of N2H+, is the Einstein coefficient of the N2H+ (10) transition, and are the statistical weights of the lower and upper energy levels, is the energy of the lower level, is the partition function estimated using the excitation temperature of the full N2H+ fits (see Eq.
+A2 from Caselli et al., 2002b).
+From the above procedure, inside the V-shaped structure, we get a total (N2H+) = 5.241015 cm-2 and a total (N2H+) = 5.910-8 M⊙, within its extent of 0.013 pc2.
+We use the average timescale of the VGs from § 5.3 (Fig.
+10), = 64.5 kyr, to determine the (N2H+) as:
+Note that the (N2H+) estimate (and (H2) below) should be multiplied by , in order to account for the unknown inclination angle () of the protocluster/filaments relative to the POS.
+For the main blue V-shape (Fig. D), we derive the N2H+ relative abundance (N2H+), using the N2H+ and H2 column densities obtained above, as:
+The N2H+) value obtained above appears lower than typical estimates in massive Galactic star forming regions reported in different works (; Caselli et al., 2002a; Henshaw et al., 2014, Sandoval-Garrido et al.
 in prep.).
+We consider it a good agreement considering the uncertainties of the involved measurements (i.e.
+column density estimates).
+For a comparison of the N2H+ relative abundance between regions composing the V-shapes and the rest of protocluster we would require to model the full N2H+ line emission in the whole region.
+Excitation temperature [K] | (Tex) | 2.73 80 | |
+Opacity | () | 0.01 40 | |
+Centroid velocity [km s-1] | () | 25 15 | |
+Velocity dispersion [km s-1] | () | 0.20 3 |
 8 Discussion
+8.1 V-shaped VGs in the literature
 The V-shaped velocity gradients described in this work have been detected across multiple Galactic star forming system.
@@
 (2021) they analyze the L1482 filament located in the California Molecular Cloud.
-In all of the analyzed tracers there is a clear velocity peak in their north region (length 1.8 pc, mass M).
+In all of the analyzed tracers there is a clear velocity peak in their north region (length 1.8 pc, mass M⊙).
 This subregion contains a higher gas density and higher number of YSOs compared to the more quiescent south part.
@@
 (2022) study the velocity profiles along filaments from the ATOMS survey (Liu et al., 2020b).
-The median mass of their sources is 1.4 M with a median length of the filaments at pc.
-By analyzing the HCO (10) emission they find converging VGs along filaments (see their Figs.
+The median mass of their sources is 1.4 M⊙ with a median length of the filaments at pc.
+By analyzing the H13CO+ (10) emission they find converging VGs along filaments (see their Figs.
 6 & 10), which they also detected using simulations from Gómez & Vázquez-Semadeni (2014).
@@
 In Zhou et al.
-(2023) they analyzed CO (21) APEX/LAsMA data of the G333 complex.
+(2023) they analyzed 13CO (21) APEX/LAsMA data of the G333 complex.
 They identify multiple V-shaped VG (see their Fig.
@@
 Redaelli et al.
-(2022) use ALMA NH (10) isolated component data of the high-mass (5200 M) clump AGAL014.492-00.139 identifying multiple coherent structures in PPV space (trees “B” and “G”; right panel of their Fig.
+(2022) use ALMA N2H+ (10) isolated component data of the high-mass (5200 M⊙) clump AGAL014.492-00.139 identifying multiple coherent structures in PPV space (trees “B” and “G”; right panel of their Fig.
 7 & 9).
@@
 For their “G” PV distribution, they suggest that one scenario is where the dense gas is flowing along the filament (of length 0.2 pc) from protostar “p3” towards the protostar “p2”.
-This motion has an = 2.2 M yr, being in the range of the we derive for our VGs (see Table D.1).
+This motion has an = 2.2 M⊙ yr-1, being in the range of the we derive for our VGs (see Table D.1).
 In Rawat et al.
-(2024) they analyze CO(10) ALMA data observing a V-shaped VG (see their Fig.
+(2024) they analyze 13CO(10) data obtained with the Purple Mountain Observatory, as part of the Milky Way Imaging Scroll Painting survey.
+They detect a V-shaped VG (see their Fig.
 14) along the ridge of the G148.24+00.41 (G148) cloud.
@@
 Note that the length of the V-shape in G148 is pc, while our most prominent V-shape (Fig.
-10) span and lengths of their identified filaments are 1.3 to 6.9 M and 14 to 38 pc respectively, large compared to the total mass (2.5 M) and extent of G353 ( pc).
-This different in probed length and masses might be reflected by their mean VG km s pc, orders of magnitude smaller than our VGs.
+10) is .
+Also the masses and lengths of their identified filaments are (1.3 to 6.9) M⊙ and 14 to 38 pc respectively, large compared to the total mass (2.5 M⊙) and extent of G353 ( pc).
+This difference in probed lengths and masses might be reflected by their mean VG km s-1 pc-1, orders of magnitude smaller than our VGs.
+This is consistent with the analysis presented in Zhou et al.
+(2022, 2023), where they observe an inverse relation between the spatial scale of a region and their velocity gradients (see their Fig.
+8).
 In Pan et al.
-(2024) using APEX CO (21) data of the filamentary cloud G034.43+00.24 (G34) they identify converging VGs of lengths pc towards the “middle ridge” see their Fig.
+(2024) using APEX C18O (21) data of the filamentary cloud G034.43+00.24 (G34) they identify converging VGs of lengths pc towards the “middle ridge” see their Fig.
 3, top panel).
 They interpret these VG as gas flowing from the filaments onto dense clumps, located at the center of G34.
-These VGs of their southern and northern filaments are in the range of km s pc, and they estimate the total mass inflow rate towards the middle ridge as M yr, similar to our estimates.
+These VGs of their southern and northern filaments are in the range of km s-1 pc-1, and they estimate the total mass inflow rate towards the middle ridge as M⊙ yr-1, similar to our estimates.
 Current work by Sandoval-Garrido et al.
@@
 (in prep) they analyze the kinematics of the evolved protocluster G012.80 (located at 2.4 kpc; Motte et al., 2022), where they implement similar techniques and find velocity signatures of filamentary rotation.
-As presented here, V-shaped VGs appear to be a generic feature across a wide range of star forming environments, probing VGs with differences of up to orders of magnitude in spatial scales ranging from 0.1 to 10 pc.
-Based on these works, the V-shaped VGs appear to be a generic feature in star forming regions.
+8.2 Filamentary 3D morphology
+V-shaped VGs appear to be a generic feature across a wide range of star forming environments, probing VGs with differences of up to orders of magnitude in spatial scales ranging from 0.1 to 10 pc.
 Despite being commonly detected in recent studies, it is still not clear how they are produced.
@@
 From our analysis regarding the most prominent V-shape (see § 5.3 & 6) we see that no core is located at its apex, although cores 2 & 3 are within pc.
-Also the position difference between the center of the V-shape with the barycenter between these two cores is AU.
+Also the spatial offset between the center of the V-shape with the barycenter between these two cores is AU.
 This is consistent with the idea of small-scale gravitational collapse within the protocluster, similar to clump decoupling from their parent molecular cloud (Peretto et al., 2023).
@@
 12).
-To shed light into this degeneracy in G353 we compare the NH (dense gas tracer) radial velocities and SiO (shock/outflow tracer) as a proxy for energies.
-The velocity range V covered by the NH emission is 8 km s, while for SiO is 80 km s.
-Given the difference in probed velocities between SiO and NH (see § 5.3) and the analysis presented in § 6 we suggest that the V-shapes in G353 are a signature of infall.
-As presented in this section, the V-shaped VGs appear to be common in recent studies of protoclusters.
-The multiple VGs that conform the V-shapes present in G353 have values of to km s pc, with timescales ranging from to kyr, and values .
+To shed light into this degeneracy in G353 we compare the N2H+ (dense gas tracer) radial velocities and SiO (shock/outflow tracer) as a proxy for energies.
+The velocity range V covered by the N2H+ emission is 8 km s-1, while for SiO is 80 km s-1.
+Given the difference in probed velocities between SiO and N2H+ (see § 5.3) and the analysis presented in § 6 we suggest that the V-shapes in G353 are a signature of infall.
+The multiple VGs that conform the V-shapes present in G353 have values of to km s-1 pc-1, with timescales ranging from to kyr, and values .
 These values are similar to VGs in other regions with comparable sizes and masses.
 In G353 it is likely that these VGs are the result of dense gas moving through filaments, possibly increasing the density of the central regions, shaping the overall velocity field at large and small scales, and leading to a further increase of the core population and star formation activity.
-One important aspect of the V-shapes that is still not well understood is the timescale associated of the VGs (VG).
+8.3 Timescales and mass accretion rates
+One important aspect of the V-shapes that is still not well understood is the timescale associated of the VGs (VG-1).
 It is not clear if nor how these timescales determine core formation lifetimes or impact the star formation environment in general.
@@
 At filament lengths comparable to our V-shapes ( pc) their gas accretion timescales are on the order of our estimates (see Table D.1).
+8.4 Depletion timescales
 It is also interesting to consider the mass accretion rates measured here compared to the available protocluster mass reservoir to explore implications for the duration of the gas dominated phase.
-The total mass accretion rate of our V-shaped structures is 310 M yr (see Table D.1).
-Considering the total mass (M) of G353 as a mass reservoir, we estimate the time needed fully consume the gas ().
+The total mass accretion rate of our V-shaped structures is 310-3 M⊙ yr-1 (see Table D.1).
+Considering the total mass (MTot) of G353 as a mass reservoir, we estimate the depletion timescale () as the time needed fully consume the gas.
 Here we assume that is representative of flows feeding gas onto cores.
-We estimate M, where the total mass of the region is 2500 M (Motte et al., 2022).
+We estimate M, where the total mass of the region is 2500 M⊙ (Motte et al., 2022).
 We obtain a = 0.8 Myr, of similar magnitude but about four times larger than the of the protocluster.
@@
 Given that the protocluster does not appear to be in a state of free-fall (see § 6) but instead undergoing comparatively slow gravitational contraction, the similarity in these relatively crude estimates seems remarkable.
-While we do not yet have an explanation for why this relatively good match in timescales, it would seem to indicate that protocluster evolution may be a self-regulating process.
+While we do not yet have an explanation for why relatively good match in timescales, it would seem to indicate that protocluster evolution may be a self-regulating process.
 Larger samples and similar analysis will test this hypothesis.
@@
 We characterize the complex dense gas kinematics of G353 using ALMA-IMF LP observations.
-The data used in this paper mainly consist of the fully combined NH data cube, but we also include 1.3 mm continuum cores and DCN cores velocity catalogues, SiO 12 m observation, and a (H) 1.3 mm continuum derived map.
+The data used in this paper mainly consist of the fully combined N2H+ data cube, but we also include 1.3 mm continuum cores and DCN cores velocity catalogues, SiO 12 m observation, and a (H2) 1.3 mm continuum derived map.
 We summarize our main results below.
 1.
-With our NH isolated component modeling, we find that most of the 1.3 mm cores are located in regions with 2 to 3 velocity components. This indicates kinematic complexity down to 4 kau scales.
+With our N2H+ isolated component modeling, we find that most of the 1.3 mm cores are located in regions with 2 to 3 velocity components. This indicates kinematic complexity down to 4 kau scales.
 2.
-We increase the number of cores with a V estimate in this region by further examining the DCN emission and comparing it with the NH data extracted towards the core positions.
+We increase the number of cores with a VLSR estimate in this region by further examining the DCN emission and comparing it with the N2H+ data extracted towards the core positions.
 We find that 11 cores, which were previously undetected in the in the DCN background-subtracted fitting from Cunningham et al.
@@
 4.
-From the PV diagrams, we see the DCN core velocities are in agreement with the NH velocity distribution (within a few DCN channel widths).
+From the PV diagrams, we see the DCN core velocities are in agreement with the N2H+ velocity distribution (within a few DCN channel widths).
 This suggests coupling between the cores and the dense gas in which they formed.
@@
 In the intensity-weighted PVs we see clear V-shaped velocity structures, composed by two linear velocity gradients (VGs) converging into a common point.
-These VGs are present across all NH velocity components.
+These VGs are present across all N2H+ velocity components.
 Some of them are near the location of cores in both position and velocity (see § 5.2)
 - 6.
-We successfully characterize nine V-shaped VGs well detected in our NH data (see Fig. D.2).
 - 7.
 8.
-We estimate the barycenter of cores 2 & 3, presenting an offset relative to the center of the V-shape of ″( 600 au) well below the beam size of our NH data.
+We estimate the barycenter of cores 2 & 3, presenting an offset relative to the center of the V-shape of ″( 600 au) well below the beam size of our N2H+ data.
 9.
-For V-shape “B” we find that core “7”, with a mass of 6 M, is located within a beam size from its apex.
+For V-shape “B” we find that core “7”, with a mass of 6 M⊙, is located within a beam size from its apex.
 10.
@@
 We characterize the VGs composing our sample of V-shapes by applying linear fits to these distributions.
-We estimate timescales associated to the VGs as VG.
+We estimate timescales associated to the VGs as VG-1.
 These timescales are between 35 to 170 kyr, with an average of 67 kyr.
-These values are short compared to the t of the protocluster ( Myr), and times larger that the cores average t ( kyr).
+These values are short compared to the tff of the protocluster ( Myr), and times larger that the cores average tff ( kyr).
 12.
-We suggest that at small scales the NH V-shaped structures indicate gas motions along filaments, towards denser regions.
+We suggest that at small scales the N2H+ V-shaped structures indicate gas motions along filaments, towards denser regions. Thus we interpret as inflow timescales.
 13.
-Using an H mass map and the V-shapes mean timescales, we derive H mass accretion rates of , consistent with previous studies on regions that present gas flows along filaments towards denser object or regions, such as protostars and clumps.
+Using an H2 mass map and the V-shapes mean timescales, we derive H2 mass accretion rates of , consistent with previous studies on regions that present gas flows along filaments towards denser object or regions, such as protostars and clumps.
+Moreover, V-shapes “H”, “C”, “F”, and “B” present the largest (H2) and they are located near or at the convergence point of the filaments (see Fig.
+13).
 14.
-In SiO, the PV structure covers a velocity range (V) of km s, while for NH V is km s.
-This difference suggests that NH is tracing infall, a less energetic processes compared to SiO, a shock and outflow tracer.
+In SiO, the PV structure covers a velocity range (V) of km s-1, while for N2H+ V is km s-1.
+This difference suggests that N2H+ is tracing infall, a less energetic processes compared to SiO, a shock and outflow tracer.
 15.
-We model a gravitationally collapsing sphere.
+We model the protocluster as a gravitationally collapsing sphere.
 The derived radial velocities are consistent with the large scale morphology of the traditional PV diagram.
@@
 Acknowledgements.
+The authors thank the referee for helpful comments that improved the text.
 This paper makes use of the following ALMA data: ADS/JAO.ALMA#2017.1.01355.L.
@@
 The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ.
-We thank Elena Redaelli and Diego R.
-Matus Carrillo for very helpful discussions.
+We thank Elena Redaelli, Diego R.
+Matus Carrillo, and Vineet Rawat for very helpful discussions.
 R.A.
@@
 is a postdoctoral fellow in the University of Virginia’s VICO collaboration and is funded by grants from the NASA Astrophysics Theory Program (grant number 80NSSC18K0558) and the NSF Astronomy & Astrophysics program (grant number 2206516).
+A.G.
+acknowledges support from the NSF under grants AAG 2008101 and CAREER 2142300.
 T.Cs.
@@
 - Fernández-López et al. (2014) Fernández-López, M., Arce, H. G., Looney, L., et al. 2014, ApJ, 790, L19
+- Galván-Madrid et al. (2013) Galván-Madrid, R., Liu, H. B., Zhang, Z. Y., et al. 2013, ApJ, 779, 121
+- Galván-Madrid et al. (2010) Galván-Madrid, R., Zhang, Q., Keto, E., et al. 2010, ApJ, 725, 17
 - Ginsburg & Mirocha (2011) Ginsburg, A. & Mirocha, J. 2011, PySpecKit: Python Spectroscopic Toolkit, Ver. 0.1.23, Astrophysics Source Code Library
@@
 - Henshaw et al. (2019) Henshaw, J. D., Ginsburg, A., Haworth, T. J., et al. 2019, MNRAS, 485, 2457
+- Henshaw et al. (2020) Henshaw, J. D., Kruijssen, J. M. D., Longmore, S. N., et al. 2020, Nature Astronomy, 4, 1064
 - Inoue & Fukui (2013) Inoue, T. & Fukui, Y. 2013, ApJ, 774, L31
@@
 - Koch & Rosolowsky (2015) Koch, E. W. & Rosolowsky, E. W. 2015, MNRAS, 452, 3435
+- Kumar et al. (2022) Kumar, M. S. N., Arzoumanian, D., Men’shchikov, A., et al. 2022, A&A, 658, A114
 - Kuznetsova et al. (2015) Kuznetsova, A., Hartmann, L., & Ballesteros-Paredes, J. 2015, ApJ, 815, 27
@@
 - Mardones et al. (1997) Mardones, D., Myers, P. C., Tafalla, M., et al. 1997, ApJ, 489, 719
+- Men’shchikov (2021) Men’shchikov, A. 2021, A&A, 649, A89
 - Minh et al. (2018) Minh, Y. C., Liu, H. B., Galvań-Madrid, R., et al. 2018, ApJ, 864, 102
@@
 - Motte et al. (2018) Motte, F., Bontemps, S., & Louvet, F. 2018, ARA&A, 56, 41
+- Myers (2009) Myers, P. C. 2009, ApJ, 700, 1609
 - Nony et al. (2023) Nony, T., Galván-Madrid, R., Motte, F., et al. 2023, A&A, 674, A75
@@
 - Pan et al. (2024) Pan, S., Liu, H.-L., & Qin, S.-L. 2024, ApJ, 960, 76
+- Peretto et al. (2014) Peretto, N., Fuller, G. A., André, P., et al. 2014, A&A, 561, A83
 - Peretto et al. (2023) Peretto, N., Rigby, A. J., Louvet, F., et al. 2023, MNRAS, 525, 2935
@@
 - Tafalla et al. (2004) Tafalla, M., Myers, P. C., Caselli, P., & Walmsley, C. M. 2004, A&A, 416, 191
+- Tafalla et al. (2021) Tafalla, M., Usero, A., & Hacar, A. 2021, A&A, 646, A97
 - Tobin et al. (2012) Tobin, J. J., Hartmann, L., Bergin, E., et al. 2012, ApJ, 748, 16
@@
 Here we describe the procedure to identify the main filamentary structures presented in § 3 using FilFinder (Koch & Rosolowsky 2015).
-For this approach we use the moment 0 map of the extracted NH isolated components that present a SNR 5.
-To estimate the moment 0 map we used the moment task from the SpectralCube Python package, within the velocity range of km s to 0 km s.
+For this approach we use the moment 0 map of the extracted N2H+ isolated components that present a SNR 5.
+To estimate the moment 0 map we used the moment task from the SpectralCube Python package, within the velocity range of km s-1 to 0 km s-1.
 As part of the pre-processing of the moment 0 map before the filamentary detection, we decrease the contrast in the image by using the preprocess_image task and its argument flatten_percent set to 90.
-Now, in order to indicate to FilFinder the area we to identify filaments use the subtask create_mask with the following parameters: glob_thresh: 4.5 K km s, size_thresh: 0.25 pc, smooth_size: 0.12 pc, border_masking: False, fill_hole_size: 0.013 pc.
+Now, in order to indicate to FilFinder the area we to identify filaments use the subtask create_mask with the following parameters: glob_thresh: 4.5 K km s-1, size_thresh: 0.25 pc2, smooth_size: 0.12 pc, border_masking: False, fill_hole_size: 0.013 pc2.
 The resulting mask is presented in Fig.
@@
 Then, we obtain the skeletons of the mask by using medskel.
-The derived structures are presented with red and black lines in Fig.
+The derived structures are presented with red and green lines in Fig.
 A.1.
@@
 This approach results in removing the small filaments (red lines in Fig.
-A.1) from the original skeleton and to obtain the main filamentary structure in G353 (black lines in Fig.
+A.1) from the original skeleton and to obtain the main filamentary structure in G353 (green lines in Fig.
 A.1).
 Appendix B Examples of the isolated components fitting
-Appendix C DCN & NH derived core velocity
+Appendix C DCN & N2H+ derived core velocity
 Appendix D V-shaped structures
@@
 In Fig.
-D.1 we indicate their location in PV space with arrows and present individual close-up for each V-shape in Fig.
-D.2.
-In Table D.1 we list their VGs, timescales, H masses, and mass accretion rates.
+D.1 we indicate the V-shapes location in PV space with dark points, arrows, and their IDs.
+In Fig.
+D.2 we present individual close-ups for each V-shape.
+In Table D.1 we list their VGs, timescales, H2 masses, and mass accretion rates.
 Here we list a few clarifications due to projection effects seen in these V-shaped structures:
@@
 For V-shape “A”, the 1.3 mm core with DCN single velocity component, located at the apex of this V-shape, is not spatially related to it.
-In V-shapes “G” and “H” we see the same 1.3 mm cores with NH velocities. These V-shapes are not the same distribution. They are overlapped in PV space and spatially separated by .
-We improve the clarity of V-shape “B” by rotating the data in PP space by 80 counter-clockwise. We apply this process for V-shapes “E”, “G”, and “H” with an angle of 33 clockwise.
+In V-shapes “G” and “H” we see the same 1.3 mm cores with N2H+ velocities. These V-shapes are not the same distribution. They are overlapped in PV space and spatially separated by .
+We improve the clarity of V-shape “B” by rotating the data in PP space by 80∘ counter-clockwise. We apply this process for V-shapes “E”, “G”, and “H” with an angle of 33∘ clockwise.
 V-shapes G and H overlap in PV space but these are structures spatially separated.
@@
 Then, we estimate the integrated intensity and velocity centroid at each pixel.
-We find improvements in our cleaning by using only spectra with integrated intensity K km s.
+We find improvements in our cleaning by using only spectra with integrated intensity K km s-1.
 Using the coordinate, integrated-intensity, and velocity centroid of each spectrum, we create the SiO intensity-weighted PV diagrams we show in E.1.
-Core | RA | DEC | F | F | PA | Mass | V | Type |
-Number | [ ] | [ ] | [″] | [″] | [ ] | [M ] | [km s] | |
-2 | 262.6165032 | -34.6955865 | 0.065 | DCN, Single | ||||
-3 | 262.6184156 | -34.6965240 | 0.040 | DCN, Single | ||||
-4 | 262.6103159 | -34.6932659 | 0.155 | DCN, Single | ||||
-5 | 262.6101515 | -34.6960014 | 0.087 | DCN, Single | ||||
-6 | 262.6049155 | -34.6934384 | 0.126 | DCN, Single | ||||
-7 | 262.6137738 | -34.6947298 | DCN & NH | |||||
-8 | 262.6039531 | -34.6936374 | ||||||
-9 | 262.6192359 | -34.6903650 | DCN & NH | |||||
-11 | 262.6243189 | -34.6880780 | DCN & NH | |||||
-12 | 262.6072148 | -34.6969795 | 0.019 | DCN, Single | ||||
-13 | 262.6078228 | -34.6996836 | ||||||
-14 | 262.6147937 | -34.6946762 | 0.090 | DCN, Single | ||||
-15 | 262.6107433 | -34.6964412 | 0.087 | DCN, Complex | ||||
-16 | 262.6215941 | -34.6989408 | ||||||
-17 | 262.5954514 | -34.6916168 | ||||||
-18 | 262.5927434 | -34.7052494 | ||||||
-19 | 262.6064012 | -34.7019756 | ||||||
-20 | 262.6111096 | -34.6932787 | ||||||
-21 | 262.6131910 | -34.6939495 | 0.063 | DCN, Single | ||||
-22 | 262.6118441 | -34.6946150 | DCN & NH | |||||
-23 | 262.6028175 | -34.6925438 | ||||||
-24 | 262.6198349 | -34.6960383 | 0.089 | DCN, Single | ||||
-25 | 262.6155222 | -34.6952591 | DCN & NH | |||||
-26 | 262.6143434 | -34.6917027 | ||||||
-27 | 262.6000802 | -34.6910324 | ||||||
-28 | 262.6253977 | -34.6999713 | ||||||
-29 | 262.6133074 | -34.6919187 | ||||||
-30 | 262.6114686 | -34.6962602 | 0.074 | DCN, Single | ||||
-31 | 262.6096651 | -34.6925680 | ||||||
-32 | 262.6094126 | -34.6910985 | ||||||
-33 | 262.6287106 | -34.6862068 | ||||||
-34 | 262.5982914 | -34.6919006 | 0.080 | DCN, Single | ||||
-35 | 262.6142011 | -34.6940134 | DCN & NH | |||||
-36 | 262.6202758 | -34.7001995 | ||||||
-37 | 262.6010398 | -34.6950114 | 0.053 | DCN, Single | ||||
-38 | 262.5971034 | -34.6920396 | ||||||
-39 | 262.6054437 | -34.6963773 | DCN & NH | |||||
-40 | 262.5917454 | -34.6897316 | ||||||
-41 | 262.6095777 | -34.6983259 | ||||||
-42 | 262.5975992 | -34.6876666 | ||||||
-43 | 262.6035166 | -34.6966807 | 0.062 | DCN, Single | ||||
-44 | 262.6030115 | -34.6956424 | 0.091 | DCN, Single | ||||
-45 | 262.6143008 | -34.6909376 | ||||||
-46 | 262.6187648 | -34.6912377 | DCN & NH | |||||
-47 | 262.6178453 | -34.6919943 |
+Core | RA | DEC | FA | FB | PA | Mass | VLSR | Type | Number of | ||
+Number | [∘ ] | [∘ ] | [″] | [″] | [∘ ] | [M⊙ ] | [km s-1] | N2H+ components | |||
+2 | 262.6165032 | -34.6955865 | 0.065 | DCN, Single | |||||||
+3 | 262.6184156 | -34.6965240 | 0.040 | DCN, Single | |||||||
+4 | 262.6103159 | -34.6932659 | 0.155 | DCN, Single | |||||||
+5 | 262.6101515 | -34.6960014 | 0.087 | DCN, Single | |||||||
+6 | 262.6049155 | -34.6934384 | 0.126 | DCN, Single | |||||||
+7 | 262.6137738 | -34.6947298 | DCN & N2H+ | ||||||||
+8 | 262.6039531 | -34.6936374 | — | — | — | ||||||
+9 | 262.6192359 | -34.6903650 | DCN & N2H+ | ||||||||
+11 | 262.6243189 | -34.6880780 | DCN & N2H+ | ||||||||
+12 | 262.6072148 | -34.6969795 | 0.019 | DCN, Single | |||||||
+13 | 262.6078228 | -34.6996836 | — | — | — | ||||||
+14 | 262.6147937 | -34.6946762 | 0.090 | DCN, Single | |||||||
+15 | 262.6107433 | -34.6964412 | 0.087 | DCN, Complex | — | ||||||
+16 | 262.6215941 | -34.6989408 | — | — | — | ||||||
+17 | 262.5954514 | -34.6916168 | — | — | — | ||||||
+18 | 262.5927434 | -34.7052494 | — | — | — | ||||||
+19 | 262.6064012 | -34.7019756 | — | — | — | ||||||
+20 | 262.6111096 | -34.6932787 | — | — | — | ||||||
+21 | 262.6131910 | -34.6939495 | 0.063 | DCN, Single | |||||||
+22 | 262.6118441 | -34.6946150 | DCN & N2H+ | ||||||||
+23 | 262.6028175 | -34.6925438 | — | — | — | ||||||
+24 | 262.6198349 | -34.6960383 | 0.089 | DCN, Single | |||||||
+25 | 262.6155222 | -34.6952591 | DCN & N2H+ | ||||||||
+26 | 262.6143434 | -34.6917027 | — | — | — | ||||||
+27 | 262.6000802 | -34.6910324 | — | — | — | ||||||
+28 | 262.6253977 | -34.6999713 | — | — | — | ||||||
+29 | 262.6133074 | -34.6919187 | — | — | — | ||||||
+30 | 262.6114686 | -34.6962602 | 0.074 | DCN, Single | |||||||
+31 | 262.6096651 | -34.6925680 | — | — | — | ||||||
+32 | 262.6094126 | -34.6910985 | — | — | — | ||||||
+33 | 262.6287106 | -34.6862068 | — | — | — | ||||||
+34 | 262.5982914 | -34.6919006 | 0.080 | DCN, Single | |||||||
+35 | 262.6142011 | -34.6940134 | DCN & N2H+ | ||||||||
+36 | 262.6202758 | -34.7001995 | — | — | — | ||||||
+37 | 262.6010398 | -34.6950114 | 0.053 | DCN, Single | |||||||
+38 | 262.5971034 | -34.6920396 | — | — | — | ||||||
+39 | 262.6054437 | -34.6963773 | DCN & N2H+ | ||||||||
+40 | 262.5917454 | -34.6897316 | — | — | — | ||||||
+41 | 262.6095777 | -34.6983259 | — | — | — | ||||||
+42 | 262.5975992 | -34.6876666 | — | — | — | ||||||
+43 | 262.6035166 | -34.6966807 | 0.062 | DCN, Single | |||||||
+44 | 262.6030115 | -34.6956424 | 0.091 | DCN, Single | |||||||
+45 | 262.6143008 | -34.6909376 | — | — | — | ||||||
+46 | 262.6187648 | -34.6912377 | DCN & N2H+ | ||||||||
+47 | 262.6178453 | -34.6919943 | — | — | — |
 […]
@@
 […]
-V-shape ID | (H) | Upper / lower VG | Mean VG | Upper / lower | (H) | |||
-[ ] | [ ] | [M] | [km s pc] | [km s pc] | [kyr] | [kyr] | [10 M yr] | |
+V-shape ID | (H2) | Upper / lower VG | Mean VG | Upper / lower | (H2) | |||
+[∘ ] | [∘ ] | [M⊙] | [km s-1 pc-1] | [km s-1 pc-1] | [kyr] | [kyr] | [10-4 M⊙ yr-1] | |
 A | 353.3981 | -0.3506 | 20.85 / 17.04 | 46.89 / 57.39 | ||||
@@
 F.2), with pc.
-Based of different testings we set the total mass of the sphere to 150 M.
-We define which corresponds to the pixel size of the NH data at a distance of 2 kpc.
-where ) = 150 M, pc, and = 5.65, from Eq. F.4, we obtain:
+Based on different tests, probing total masses from M⊙ and values from , we set the total mass of the sphere to 150 M⊙.
+We define which corresponds to the pixel size of the N2H+ data at a distance of 2 kpc.
+where ) = 150 M⊙, pc, and = 5.65, from Eq. F.4, we obtain: